### v2

# Relative Entropy Methods for the Approximation of Reactive Trajectories thanks: BvK was supported by NSF DMS-2012207

Abstract
Motivated by challenges arising in molecular simulation, we study reactive trajectories of the overdamped Langevin dynamics, i.e. trajectories observed as they pass from a set $A$ corresponding to the reagents of a chemical reaction to a set $B$ corresponding to the products. Reactive trajectories are known to have the same distribution as trajectories of the overdamped Langevin dynamics biased by a singular drift related to the committor function. In this work, we assess the effect of replacing the exact singular drift with an approximation based on an approximate committor function. We derive a convenient formula for the relative entropy between the distributions of exact and approximate reactive trajectories, and we propose a stochastic gradient descent method for minimizing the entropy to train an approximate committor function on the fly while computing reactive trajectories. We also devise a model assessment procedure for comparing the qualities of different approximations to the committor function based on the relative entropy.
1 Introduction
For many molecular systems, the most interesting phenomena relate to extremely rare transitions between long-lived states, e.g. transitions between folded states of a protein. Such systems are called metastable. It may be infeasible to perform a simulation long enough to observe even one rare transition in a metastable system. Therefore, a wide variety of methods have been devised to compute statistics of rare transitions by means other than direct simulation. Some of these methods, like adaptive multilevel splitting [9, 10], forward flux sampling [1], and transition path sampling [6] produce samples of transition paths or fragments of transition paths. Others produce a representative path, for example the minimizer of the Wentzell–Friedlin action, which is roughly speaking the mode of the distribution of transition paths in the limit of low temperature [29]; see [13] for a survey. Finally, one can compute many important statistics of transition paths in terms of solutions of partial differential equations using transition path theory [12, 13]. By various means, these methods efficiently explore regions associated with rare transitions.
Motivated by the challenge of simulating metastable systems, we study reactive trajectories of the overdamped Langevin dynamics
(1) $\mathrm{d}X_{t}=-\nabla U(X_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t},$
which are trajectories observed as they pass from a set $A$ corresponding to the reagents of a chemical reaction to a set $B$ corresponding to the products; cf. Figures 1 and 2. In [24], building on the transition path theory of E and Vanden-Eijnden [12], Lu and Nolen showed that reactive trajectories have the same distribution as the unique strong solution of the transition path equation
(2) $\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
with initial condition
(3) $Y_{0}\sim\frac{1}{\zeta}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x).$
Here,
$q(x)=\mathds{P}[\tau_{B}<\tau_{A}|Y_{0}=x]$
is the committor function, i.e. the probability that a trajectory of the overdamped Langevin dynamics (1) with $Y_{0}=x$ will hit $B$ before $A$, $S_{A}$ is the surface measure on $\partial A$, and $\zeta$ is a normalizing constant. The drift term $2\varepsilon\nabla\log q(X_{t})$ in (2) is singular on $\partial A$, since $q$ is zero on $\partial A$, and it forces trajectories away from $A$ and towards $B$ so that the lengths of reactive trajectories will typically be much shorter than the intervals of time between reactions; cf. Figure 2. That is, the singular drift is a bias that forces rare transitions to occur quickly and according to the correct distribution.
[…]
To explain why we are interested in methods of this type, we wish to make two comparisons with other classes of methods in the computational chemistry literature. First, adaptive multilevel splitting [9] computes a sample from the exact distribution of reactive trajectories initiated from a fixed starting point near the boundary of $A$ by branching and killing trajectory fragments. Our methods produce approximate reactive trajectories, but since we calculate the committor and trajectories together, we can ensure that the initial points have at least approximately the reactive flux distribution (3). [^2] We note that, in [23], Lopes and Lelièvre propose a very promising means of determining the right distribution of initial points for adaptive multilevel splitting, but their method requires the choice of a surface close to $A$ having certain properties that might be difficult to verify in practice. Our approach does not require the identification of any such surface. Moreover, unlike adaptive multilevel splitting, our approach does not require the explicit specification of a reaction coordinate. In fact, while computing reactive trajectories, we solve for the committor, which is in some respects the ideal reaction coordinate [10].
[…]
To summarize, our proposed relative entropy optimization method computes both reactive trajectories and the committor, with error estimates, with at least approximately the correct distribution of initial points, and without requiring an explicit choice of collocation points or a reaction coordinate. Instead, the user must specify an initial approximation $\tilde{q}_{0}$ to the committor function. We discuss the choice of $\tilde{q}_{0}$ in Section 6. Here, we note only that one could choose $\tilde{q}_{0}$ so that the transition path process with $\tilde{q}_{0}$ in place of $q$ imposes a biasing force pushing on a small number of atoms to induce transitions to occur as in steered molecular dynamics [17].
In addition to optimizing the relative entropy, we demonstrate two other potential applications of our theoretical results: The assessment of models of the committor and importance sampling. For high-dimensional systems, the committor and related quantities such as the density and current of transition paths may be difficult to interpret without coarse-graining, i.e. dimensionality reduction. We suggest that one could select or optimize coarse models of the committor based on the relative entropy; cf. Section 6. We also note that one can use importance sampling to correct for errors in the distribution of reactive trajectories generated based on an approximate committor. However, the change of measure involves an exponential average, and experience with similar importance sampling methods based on Girsanov’s theorem suggests that such an approach would not be reliable for poor approximations of the committor; cf. Section 4.4. Nonetheless, importance sampling yields good results for our simple test problem in Section 6.5.
We warn the reader that all theoretical and computational applications of our results demand that care be taken to correctly handle singularities. For example, since the drift is singular standard numerical integrators may fail. We have proposed an operator splitting scheme for integrating (2) in the case where $\partial A$ is planar. We do not prove convergence of the integrator, nor do we propose an integrator that works when $\partial A$ is curved, although we do grant that one would often want a curved boundary. We also note that the change of measure formula (12) involves an integral whose integrand may be singular; cf. Section 5.2.
Finally, we note that we are only able to prove the existence of the gradient $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{q_{\theta}}\|\mathds{Q})$ under rather stringent conditions; cf. Assumption 4.3. We suspect, but cannot prove, that the relative entropy is differentiable under weaker conditions. In any case, our method of training the committor by stochastic gradient descent appears to work well even when Assumption 4.3 fails; cf. Section 6.4. Note that we require Assumption 4.3 only to prove differentiability, not to derive the change of measure formula in Theorem 3.11 or to prove the weak law of large numbers for the estimator of relative entropy differences in Section 4.2.
Outline. In Section 2, we briefly review transition path theory. In Section 3, we derive the change of measure between the exact distribution of transition paths and the distribution of (2) with an approximate committor $\tilde{q}$ in place of $q$. In Section 4, we explain how to estimate relative entropy differences and gradients, and we consider the possibility of importance sampling. In Section 5, we address some practical obstacles to the implementation of our methods, including strategies for the representation of approximate committor functions that satisfy Assumption 3 and Assumption 4.3. Finally, in Section 6, we apply our methods to a two-dimensional toy model. First, we propose an integrator capable of handling the singular drift. We then assess the quality of a coarse approximation of a committor, train an improved approximation by stochastic gradient descent, demonstrate the use of importance sampling to correct for errors in the committor, and study the convergence of the integrator.
2 A Digest of Transition Path Theory
In this section, we present a brief review of transition path theory [12, 24]. Let $U:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a potential energy, describing for example a molecular system. Let $\varepsilon=k_{B}T>0$ be a temperature parameter, and suppose that $X_{t}$ evolves according to the overdamped Langevin dynamics
(4) $\mathrm{d}X_{t}=-\nabla U(X_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}.$
Under certain conditions on $U$ [20], the overdamped Langevin dynamics is ergodic for the Boltzmann distribution
$\rho(\mathrm{d}x)=Z^{-1}\exp(-U(x)/\varepsilon)\,\mathrm{d}x\text{ where }Z=\int_{\mathbb{R}^{d}}\exp(-U(x)/\varepsilon)\,\mathrm{d}x.$
We assume ergodicity throughout this work.
{assumption}
For the given potential function $U$, the overdamped Langevin dynamics (4) is ergodic.
In transition path theory, to characterize rare transitions of the overdamped Langevin dynamics, one first chooses a disjoint pair of subsets $A$ and $B$ of $\mathbb{R}^{d}$. In applications, $A$ and $B$ would usually be associated with metastable states, e.g. they might be sets modeling different conformations of a biomolecule or the reagents and products of a chemical reaction. In addition to the sets $A$ and $B$, we find it convenient to define the transition region
$\mathscr{T}=\mathbb{R}^{d}\setminus(A\cup B).$
We impose the following assumptions on $A$, $B$, and $\mathscr{T}$.
{assumption}
We assume that $A$ and $B$ are disjoint, closed subsets of $\mathbb{R}^{d}$ with nonempty interior. The boundaries of $A$ and $B$ are regular $C^{\infty}$-submanifolds of $\mathbb{R}^{d}$. The transition region $\mathscr{T}$ is connected.
Figure 1: Schematic illustration of reactive trajectories, entrance times, and exit times. The orange line depicts a trajectory of the overdamped Langevin dynamics. The bold red segments depict reactive trajectories. Here, $\tau_{B,0}$ is the first hitting time of $B$, $\tau_{A,1}$ is the first hitting time of AA after $\tau_{B,0}$, $\tau_{B,1}$ is the first hitting time of $B$ after $\tau_{A,1}$, and so forth. The $k$’th exit time $\sigma_{A,k}$ is the last time before $\tau_{B,k}$ that the process was in $A$.
Figure 2: Reactive trajectories. On the left, we have a single long trajectory of the overdamped Langevin dynamics for the simple two-dimensional model potential introduced in Section 6. The blue curves are contours of the potential. The trajectory was initialized from the Boltzmann distribution. Its starting point happened to be in the set AA on the left side of the figure. The trajectory was terminated on hitting the boundary of $B$ for the first time. The bold red curve depicts the end of the trajectory from the last time it left AA to the first time it entered $B$, i.e. the reactive part of the trajectory. The faint orange curve depicts the beginning of the trajectory up to the last time it left AA. On the right, we have three approximate reactive trajectories computed by integrating the transition path process for the approximate committor $q_{\rm{sgd}}$ calculated by gradient descent in Section 6.4. The blue curves are contours of the effective potential $U-2\varepsilon\log q_{\rm{sgd}}$ for the transition path process, which is singular on $\partial A$. The red and pink curves are the reactive trajectories.
Given a pair of sets $A$ and $B$, one defines sequences of entrance times, exit times, and reactive trajectories as illustrated in Figure 1. The entrance times are
$\tau_{A,0}$ $:=\inf\{t\geq 0;X_{t}\in A\},$
$\tau_{B,0}$ $:=\inf\{t\geq\tau_{A,0};X_{t}\in B\},$
and for $k\geq 1$
$\tau_{A,k}$ $:=\inf\{t\geq\tau_{B,k-1};X_{t}\in A\},$
$\tau_{B,k}$ $:=\inf\{t\geq\tau_{A,k-1};X_{t}\in B\}.$
The exit times are
$\sigma_{A,0}$ $:=\sup\{\tau_{A_{0}}\leq t\leq\tau_{B,0};X_{t}\in A\},$
$\sigma_{B,0}$ $:=\sup\{\tau_{B,0}\leq t\leq\tau_{A,1};X_{t}\in B\},$
and for $k\geq 1$
$\sigma_{A,k}$ $:=\sup\{\tau_{A,k}\leq t\leq\tau_{B,k};X_{t}\in A\},$
$\sigma_{B,k}$ $:=\sup\{\tau_{B,0}\leq t\leq\tau_{A,1};X_{t}\in B\}.$
The $k$’th reactive trajectory (or transition path) $Y^{k}_{t}$ is the segment of the random path $X_{t}$ between the $k$’th exit time $\sigma_{A,k}$ from $A$ and the $k$’th entrance time $\tau_{B,k}$ to $B$. That is,
$Y^{k}_{t}:=X_{\sigma_{A,k}+t}\text{ for }t\in[0,\tau_{B,k}-\sigma_{A,k}].$
Transition path theory characterizes the distribution of reactive trajectories in terms of the committor function
$q(x):=\mathds{P}[\tau_{B}<\tau_{A}|X_{0}=x].$
Here, $\tau_{A}$ and $\tau_{B}$ are the first hitting times of $A$ and $B$ for $X_{t}$, so $q(x)$ is the probability of hitting $B$ before $A$ when starting from $x$. In [24], Lu and Nolen verified that if the overdamped Langevin dynamics is in the steady state with $X_{t}\sim\rho$ for all $t\geq 0$, then the reactive trajectories have the same distribution as the solution of the transition path equation
(5) $\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
where $Y_{0}$ has the reactive flux distribution
(6) $Y_{0}\sim\frac{1}{\zeta}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S(x).$
Here, $S$ is the surface measure on $\partial A$, and
$\zeta=\int_{\partial A}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S(x).$
Observe that the transition path equation is the overdamped Langevin dynamics with the additional drift term $2\varepsilon\nabla\log q(X_{t})\,\mathrm{d}t$. On $\partial A$, we have $q=0$, and by Lemma 2.2 below, $\lvert\nabla q\rvert>0$. Therefore, the drift $\nabla\log q=\frac{\nabla q}{q}$ is singular on $\partial A$. Despite the singularity, Lu and Nolen showed that a unique strong solution of (5) exists for any initial distribution, even distributions supported on $\partial A$. Moreover, with probability one, $X_{t}\notin A$ for $t>0$. In effect, the singular drift repels trajectories from $A$, forcing transitions to occur.
Remark 2.1.
One can derive (5) formally by conditioning the overdamped Langevin dynamics on hitting $B$ before $A$ using the Doob $h$-transform. This approach does not establish a definite connection with reactive trajectories or the existence of strong solutions.
The committor solves the elliptic boundary value problem
(7) $\begin{cases}Lq=0&\text{ in }\mathscr{T},\\ q=0&\text{ on }\partial A,\\ q=1&\text{ on }\partial B,\end{cases}$
where
$L:=-\nabla U\cdot\nabla+\varepsilon\Delta$
is the generator of the overdamped Langevin dynamics [12]. Therefore, under our smoothness assumptions on $U$ and $\mathscr{T}$, elliptic regularity, the Hopf lemma, and the maximum principle imply the properties of the committor listed in Lemma 2.2 below.
{assumption}
We assume that the potential energy $U:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is infinitely differentiable, but we do not assume that $U$ or any of its derivatives are bounded.
Lemma 2.2.
Under Assumptions 2 and 2, the forward committor function $q:\mathscr{T}\rightarrow[0,1]$ extends to an infinitely differentiable function defined on an open set containing $\partial A$ and $\partial B$. By abuse of notation, we let $q$ refer to both the committor function and the extension. For all $x\in\mathring{\mathscr{T}}$,
$q(x)>0.$
For all $x\in\partial A$,
$\nabla q(x)\cdot n(x)=\lvert\nabla q(x)\rvert>0,$
where $n(x)$ is the outward unit normal to $A$ at $x$.
Proof 2.3.
By elliptic regularity, there exists an infinitely differentiable extension of $q$ to an open set containing $\overline{\mathscr{T}}$. We have $\nabla q(x)\cdot n(x)>0$ for $x\in\partial A$ by the Hopf Lemma, and $q(x)>0$ in the interior of $\mathscr{T}$ by the maximum principle. Moreover, since $q$ is constant on $\partial A$ and $\nabla q(x)\cdot n(x)>0$, $\nabla q(x)=\lvert\nabla q(x)\rvert n(x)$.
Remark 2.4.
It is conventional to take the domain of the committor to be $\mathbb{R}^{d}$ instead of $\mathscr{T}$, letting $q=0$ on $A$ and $q=1$ on $B$. This extension of $q$ does not coincide with the smooth extension described in Lemma 2.2. Any smooth extension must take negative values at some points in the interior of $A$ since $\nabla q\cdot n>0$ and $q=0$ on $\partial A$. Moreover, observe that for our smooth extension, $Lq=0$ everywhere in $\bar{\mathscr{T}}$, including on $\partial A$. In the original work on transition state theory, $Lq$ was instead understood as a distribution supported on $\partial A\cup\partial B$. We require a smooth extension only to simplify the notation in certain proofs. None of our results depends on a particular choice of extension.
3 A Change of Measure Formula for Approximations to the Transition Path Process
We analyze the errors that result when one substitutes an approximation $\tilde{q}$ of the committor in the transition path equation (5). To be precise, in Theorem 3.11 below, we present a change of measure formula relating the distribution of solutions of the approximate transition path equation
(8) $\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log\tilde{q}(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
with the exact transition path process. To derive the change of measure, motivated by Lemma 2.2, we impose the following assumptions on the approximate committor $\tilde{q}$.
{assumption}
Let $\tilde{q}:\mathscr{T}\rightarrow\mathbb{R}$ be an approximation to the committor with the following properties:
  1. 1.
    $\tilde{q}$ extends to an infinitely differentiable function defined on an open set containing $\partial A$ and $\partial B$.
  2. 2.
    $\tilde{q}(x)=0$ for all $x\in\partial A$.
  3. 3.
    $\nabla\tilde{q}(x)\cdot n(x)>0$ for all $x\in\partial A$ where $n(x)$ is the outward unit normal to $A$ at $x$.
We now show that when an approximate committor $\tilde{q}$ has the properties outlined above, the ratio $\frac{\tilde{q}}{q}$ must be smooth in a neighborhood of $\partial A$. This will be the crucial property in deriving sufficient conditions for a change of measure.
Lemma 3.1.
Let $q$ be the forward committor. Let $\tilde{q}$ be an approximation to $q$ having the properties outlined in Assumption 3. The function
$r(x)=\frac{\tilde{q}(x)}{q(x)}$
defined for $x\in\mathring{\mathscr{T}}$ extends to a function in $C^{\infty}(\overline{\mathscr{T}})$. By abuse of notation, we let $r$ denote both the function and its extension. For $x\in\partial A$, we have
$r(x)=\frac{\lvert\nabla\tilde{q}(x)\rvert}{\lvert\nabla q(x)\rvert}=\frac{\nabla\tilde{q}(x)\cdot n(x)}{\nabla q(x)\cdot n(x)}.$
Proof 3.2.
Under Assumption 2, there exists a tubular neighborhood of $\partial A$. That is, there exist an open set $\mathscr{D}$ containing $\partial A$, an infinitely differentiable retraction $\rho:\mathscr{D}\rightarrow\partial A$, and an infinitely differentiable mapping $h:\mathscr{D}\rightarrow\mathbb{R}$ so that for any $x\in\mathscr{D}$,
$x=\rho(x)+h(x)n(\rho(x))$
where $n(y)$ is the outward unit normal to $A$ at $y\in\partial A$. To simplify notation, we define
$\bar{x}:=\rho(x),$
and we sometimes write $h$ and $n$ for $h(x)$ and $n(\rho(x))$. We choose $\mathscr{D}$ so that $\mathscr{D}$ and $B$ are disjoint.
Since $\tilde{q}>0$ on $\mathscr{T}\cup\partial B$, $r$ is smooth in an open neighborhood of any point in $\mathscr{T}\cup\partial B$. It will suffice to show that $r$ is smooth up to $\partial A$. Let $x\in\mathscr{D}$. By Taylor’s theorem and Lemma 2.2,
$q(x)$ $=q(\bar{x}+hn)$
$=q(\bar{x})+(\nabla q(\bar{x})\cdot n)h+h^{2}\int_{s=0}^{1}n^{t}D^{2}q(\bar{x}+shn)n(1-s)\,\mathrm{d}s$
$=\lvert\nabla q(\bar{x})\rvert h+h^{2}\int_{s=0}^{1}n^{t}D^{2}q(\bar{x}+shn)n(1-s)\,\mathrm{d}s.$
Similarly, we have
$\tilde{q}(x)=\lvert\nabla\tilde{q}(\bar{x})\rvert h+h^{2}\int_{s=0}^{1}n^{t}D^{2}\tilde{q}(\bar{x}+shn)n(1-s)\,\mathrm{d}s,$
since the proof of Lemma 2.2 verifies that $\nabla\tilde{q}(\bar{x})=\lvert\nabla\tilde{q}(\bar{x})\rvert n$ under our assumptions on $\tilde{q}$. Therefore,
$\frac{\tilde{q}(x)}{q(x)}=\frac{\lvert\nabla\tilde{q}(\bar{x})\rvert}{\lvert\nabla q(\bar{x})\rvert}\frac{1+\frac{h}{\lvert\nabla\tilde{q}(\bar{x})\rvert}\int_{s=0}^{1}n^{t}D^{2}\tilde{q}(\bar{x}+shn)n(1-s)\,\mathrm{d}s}{1+\frac{h}{\lvert\nabla q(\bar{x})\rvert}\int_{s=0}^{1}n^{t}D^{2}q(\bar{x}+shn)n(1-s)\,\mathrm{d}s}.$
The result follows since $\bar{x}$ and $h$ are smooth functions of $x$, $\tilde{q}$ and $q$ are in $C^{\infty}(\overline{\mathscr{T}})$, and $\lvert\nabla\tilde{q}(\bar{x})\rvert$ and $\lvert\nabla\tilde{q}(\bar{x})\rvert$ are positive.
Lemma 3.1 implies that $\log r$ and $\nabla\log r$ are smooth and bounded in a neighborhood of any point on $\partial A$, since $\lvert\nabla q(x)\rvert>0$ for all $x\in\partial A$ by Lemma 2.2. If $\nabla\log r$ is in fact globally bounded, then the Novikov condition stated in Assumption 3 below holds, and Girsanov’s theorem guarantees a change of measure.
{assumption}
For any $t>0$, we have the Novikov condition
$\mathds{E}_{\mathds{Q}}\left[\exp\left(\varepsilon\int_{0}^{t}\lvert\nabla\log r(X_{s})\rvert^{2}\,\mathrm{d}s\right)\right]<\infty.$
Establishing or assuming a global bound on $\nabla\log r$ seems to us to be the most practical means of verifying the Novikov condition and the existence of a change of measure. For example, a global bound on $\nabla\log r$ follows from Assumption 3 and Lemma 3.1 when $\mathscr{T}$ is bounded. We note, however, that a global bound is only a convenient sufficient condition, not necessary, for the results below.
To state our change of measure, we require some additional notation to properly specify the probability space on which the exact and approximate transition path processes are defined. Let $B_{t}$ be a $d$-dimensional standard Brownian motion on the probability space $(\Omega,\mathcal{F},\mathds{Q})$, and suppose that $Y_{0}$ is a random variable on $(\Omega,\mathcal{F})$ taking values in $\overline{\mathscr{T}}$ and independent of $B_{t}$. For now, we do not make any assumptions regarding the distribution of $Y_{0}$, but, in Theorem 3.11 below, $Y_{0}$ will have the reactive flux distribution (6). Let $\mathcal{F}_{t}$ be the filtration generated by $B_{t}$ and $Y_{0}$. In [24], it is shown that a strong solution of
(9) $dY_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
exists with initial value $Y_{0}$. Under our assumptions, Girsanov’s theorem yields a change of measure relating the transition path process (9) to a weak solution of the approximate transition path equation
$\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log\tilde{q}(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}.$
Lemma 3.3.
Let Assumptions 223, and 3 hold. Let $Y_{t}$ and $Y_{0}$ be as described in the previous paragraph. The process
(10) $M_{t}:=\exp\left(\int_{0}^{t}\sqrt{2\varepsilon}\nabla\log r(Y_{s})\cdot\mathrm{d}B_{s}-\frac{1}{2}\int_{0}^{t}\lvert\sqrt{2\varepsilon}\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right)$
is a positive $\mathcal{F}_{t}$-martingale under $\mathds{Q}$ with $\mathds{E}_{\mathds{Q}}[M_{t}]=1$ for all $t\geq 0$. Therefore, there exists a probability measure $\mathds{P}$ on $(\Omega,\mathcal{F})$ so that for any fixed $T>0$,
$\left.\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}\right\rvert_{\mathcal{F}_{T}}=M_{T}.$
Define the process $W_{t}$ by
$\sqrt{2\varepsilon}W_{t}:=Y_{t}-Y_{0}-\int_{0}^{t}-\nabla U(Y_{s})+2\varepsilon\nabla\log\tilde{q}(Y_{s})\,\mathrm{d}s$
so that $Y_{t}$ solves
$\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log\tilde{q}(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}W_{t}.$
Under $\mathds{P}$, $\{W_{t}\}_{t=0}^{T}$ has the law of a standard Brownian motion.
Proof 3.4.
The above merely restates Girsanov’s theorem; cf.[25, Section 8.6]. Observe that Girsanov’s theorem holds even when the drift is singular. Only some form of Novikov’s condition is required.
Remark 3.5 (Weak Solutions of the Approximate Transition Path Equation).
Lemma 3.3 establishes the existence of at least a weak solution to
$\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log\tilde{q}(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
with arbitrary initial distribution supported on $\partial A$ even though $\nabla\log\tilde{q}$ is singular on $\partial A$. We do not verify the existence of strong solutions. However, we suspect that one could do so based on arguments roughly similar to those used in [24] to construct strong solutions of the transition path equation (5).
We now derive a formula for the change of measure that can be computed in practice. Observe that formula (10) for $M_{t}$ cannot be computed without knowledge of the exact committor $q$. As a first step towards eliminating this dependence on the committor, we show that $\log q$ solves a Hamilton–Jacobi–Bellman equation. Results of this nature are standard; we include a proof only for the reader’s convenience.
Lemma 3.6.
For $x\in\mathring{\mathscr{T}}$,
$0$ $=-\nabla U(x)\cdot\nabla\log q(x)+\varepsilon\Delta\log q(x)+\varepsilon\lvert\nabla\log q(x)\rvert^{2}$
$=L\log q(x)+\varepsilon\lvert\nabla\log q(x)\rvert^{2}.$
Proof 3.7.
Define
$m(x):=\log q(x)$
for $x\in\mathscr{T}$. We have
$0$ $=Lq$
$=-\nabla U\cdot\nabla q+\varepsilon\Delta q$
$=-\nabla U\cdot\nabla\exp(m)+\varepsilon\Delta\exp(m)$
$=\exp(m)(-\nabla U\cdot\nabla m+\varepsilon\Delta m+\varepsilon\lvert\nabla m\rvert^{2})$
on $x\in\mathbb{R}^{d}\setminus(A\cup B)$, which implies the result since $\exp(m)$ must be positive.
We use the Hamilton–Jacobi–Bellman equation in Lemma 3.6 to eliminate $q$ from the exponential factor in formula (10) for $M_{t}$.
Lemma 3.8.
Under the hypotheses of Lemma 3.3, and assuming $Y_{0}$ takes values in $\partial A$, we have
$M_{t}$ $=\frac{\lvert\nabla q(Y_{0})\rvert}{\lvert\nabla\tilde{q}(Y_{0})\rvert}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}\exp\left(-\int_{0}^{t}L\log\tilde{q}(Y_{s})+\varepsilon\lvert\nabla\log\tilde{q}(Y_{s})\rvert^{2}\,\mathrm{d}s\right)$
(11) $=\frac{\lvert\nabla q(Y_{0})\rvert}{\lvert\nabla\tilde{q}(Y_{0})\rvert}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}\exp\left(-\int_{0}^{t}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)$
for $t>0$ and $M_{0}=1$.
Proof 3.9.
First, we derive an alternative expression for the integral with respect to $\mathrm{d}B_{s}$ that appears in formula (10) for $M_{t}$. By the Ito formula, for any $g\in C^{2}(\mathbb{R}^{d};\mathbb{R})$, we have
$\mathrm{d}g(Y_{t})$ $=\nabla g(Y_{t})^{t}\mathrm{d}Y_{t}+\frac{1}{2}\mathrm{d}Y_{t}^{t}D^{2}g(Y_{t})\mathrm{d}Y_{t}$
$=Lg(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})^{t}\nabla g(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\nabla g(Y_{t})^{t}\mathrm{d}B_{t},$
so
$\sqrt{2\varepsilon}\int_{0}^{t}\nabla g(Y_{s})^{t}\mathrm{d}B_{s}=g(Y_{t})-g(Y_{0})-\int_{0}^{t}Lg(Y_{s})+2\varepsilon\nabla\log q(Y_{s})^{t}\nabla g(Y_{s})\,\mathrm{d}s.$
Applying the above with $g=\log r$ and using Lemma 3.1, we have for $t>0$ that
$M_{t}$ $=\frac{r(Y_{t})}{r(Y_{0})}\exp\left(-\int_{0}^{t}L\log r(Y_{s})+2\varepsilon\nabla\log q(Y_{s})^{t}\nabla\log r(Y_{s})+\varepsilon\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right)$
$=\frac{\lvert\nabla q(Y_{0})\rvert}{\lvert\tilde{\nabla}q(Y_{0})\rvert}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}$
$\quad\times\exp\left(-\int_{0}^{t}L\log r(Y_{s})+2\varepsilon\nabla\log q(Y_{s})^{t}\nabla\log r(Y_{s})+\varepsilon\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right).$
By Lemma 3.6,
$L\log r+2\varepsilon\nabla\log q\cdot\nabla\log r+\varepsilon\lvert\nabla\log r\rvert^{2}$
$\qquad=L\log\tilde{q}-L\log q+2\varepsilon\nabla\log q\cdot\nabla\log\tilde{q}-2\varepsilon\lvert\nabla\log q\rvert^{2}$
$\qquad\qquad+\varepsilon\lvert\nabla\log\tilde{q}\rvert^{2}+\varepsilon\lvert\nabla\log q\rvert^{2}-2\varepsilon\nabla\log q\cdot\nabla\log\tilde{q}$
$\qquad=L\log\tilde{q}+\frac{1}{2}\lvert\nabla\log\tilde{q}\rvert^{2}.$
Thus,
$M_{t}=\frac{\lvert\nabla q(Y_{0})\rvert}{\lvert\tilde{\nabla}q(Y_{0})\rvert}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}\exp\left(-\int_{0}^{t}L\log\tilde{q}(Y_{s})+\varepsilon\lvert\nabla\log\tilde{q}(Y_{s})\rvert^{2}\,\mathrm{d}s\right).$
Here, for $t>0$, we have $Y_{t}\in\mathring{\mathscr{T}}$ with probability one, so $q(Y_{t})>0$, and the right-hand-side of the equation above is well-defined. In addition, $M_{0}=1$ by (10). This verifies the first formula for $M_{t}$ in the statement of the lemma. The second formula follows from the first and the identity
$L\log\tilde{q}$ $=\varepsilon\Delta\log\tilde{q}-\nabla U\cdot\nabla\log\tilde{q}$
$=\frac{L\tilde{q}}{\tilde{q}}-\varepsilon\lvert\nabla\log\tilde{q}\rvert^{2}.$
We must now eliminate the factors $\lvert\nabla q(Y_{0})\rvert$ and $q(Y_{t})$ from the change of measure formula (11). To get rid of $q(Y_{t})$, we simply stop observing the process at the first hitting time $\tau$ of $B$; note that $Y_{\tau}\in\partial B$, so $q(Y_{\tau})=1$. Observing the process up to time $\tau$ corresponds to restricting the measures $\mathds{P}$ and $\mathds{Q}$ to the stopping time $\sigma$-algebra defined below.
Definition 3.10.
Let $\tau$ be the first hitting time of $B$ for the process $Y_{t}$ solving (9). Let $\mathcal{F}_{\tau}$ be the stopping time $\sigma$-algebra of $\tau$, i.e.
$\mathcal{F}_{\tau}:=\{A\in\mathcal{F}:A\cap\{\tau\leq t\}\in\mathcal{F}_{t}\text{ for all }t\geq 0\}.$
For Theorem 3.11 below to hold, $\tau$ must be finite for both the exact and approximate transition path processes. Assumption 2 on the ergodicity of the overdamped Langevin dynamics implies that the exact transition path process will hit $B$ with probability one, since $B$ has positive Lebesgue measure. That is, $\mathds{Q}[\tau<\infty]=1$. We assume that $\tau$ is finite for the approximate process as well.
{assumption}
We assume that
$\mathds{P}[\tau<\infty]=1$
for any initial value $Y_{0}$ supported in $\overline{\mathscr{T}}$. Recall that $\mathds{P}$ is the measure introduced in Lemma 3.3 under which $Y_{t}$ is a weak solution of the approximate transition path equation (8).
We eliminate $\lvert\nabla q(Y_{0})\rvert$ from (11) by taking $Y_{0}$ to have the reactive flux distribution under $\mathds{Q}$. In that case, the factor $\lvert\nabla q(Y_{0})\rvert$ in the reactive flux (6) cancels with the one in (11). Everything that remains in the change of measure can be computed without knowing the exact committor $q$, except for a normalizing constant. The result is formula (12) in Theorem 3.11.
Theorem 3.11.
Let Assumptions 23 hold. Let $B_{t}$ be a $d$-dimensional standard Brownian motion on the probability space $(\Omega,\mathcal{F},\mathds{Q})$. Let $Y_{0}$ be a random variable on $(\Omega,\mathcal{F})$ that has the reactive flux distribution
$Y_{0}\sim\frac{1}{\zeta}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x),$
and assume that $Y_{0}$ is independent of $B_{t}$ under $\mathds{Q}$. Let $\mathcal{F}_{t}$ be the filtration generated by $B_{t}$ and $Y_{0}$. Suppose that $Y_{t}$ solves the transition path equation
$\mathrm{d}Y_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
with initial condition $Y_{0}$. Now let $\tilde{q}$ be an approximation to the committor $q$, and let $m:\partial A\rightarrow[0,\infty)$ be an unnormalized density over $\partial A$ that approximates the reactive flux distribution. Let
$\eta=\int_{\partial A}m(x)\,\mathrm{d}S_{A}(x)$
be the normalizing constant of $m$. For $t>0$, define
$Z_{t}:=\frac{\zeta}{\eta}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}\frac{m(Y_{0})}{\lvert\nabla\tilde{q}(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\exp\left(-\int_{0}^{t}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right),$
and let
$Z_{0}:=\frac{\zeta}{\eta}\frac{m(Y_{0})}{\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}.$
The process $Z_{t}$ is a nonnegative $\mathcal{F}_{t}$-martingale with $\mathds{E}_{\mathds{Q}}[Z_{t}]=1$ for all $t\geq 0$, so there exists a probability measure $\mathds{P}$ on $(\Omega,\mathcal{F})$ with
$\left.\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}\right\rvert_{\mathcal{F}_{t}}=Z_{t}$
for all $t\geq 0$. Under $\mathds{P}$,
$W_{t}:=Y_{t}-Y_{0}-\int_{0}^{t}-\nabla U(Y_{s})+2\varepsilon\nabla\log\tilde{q}(Y_{s})\,\mathrm{d}s$
is a Brownian motion and $Y_{0}\sim\eta^{-1}m(x)\,\mathrm{d}S_{A}(x)$. The density of $\mathds{P}$ restricted to the stopping time $\sigma$-algebra $\mathcal{F}_{\tau}$ is
(12) $\left.\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}\right\rvert_{\mathcal{F}_{\tau}}=Z_{\tau}=\frac{\zeta}{\eta}\tilde{q}(Y_{\tau})\frac{m(Y_{0})}{\lvert\nabla\tilde{q}(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\exp\left(-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right).$
Proof 3.12.
See Appendix A.
4 Applications of the Change of Measure
In this section, we consider two applications of the change of measure formula: selection and training. By training, we mean the minimization of the relative entropy $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ of the approximate transition path process with respect to the exact process over a family of approximate committor functions. By selection, we mean the comparison of different approximations to the committor function based on relative entropy differences. This is somewhat similar to model selection based on the Bayes and Akaike information criteria except that we do not consider penalizing model complexity. We also speculate on two other possibilities: error estimation and importance sampling. By error estimation, we mean the direct estimation of $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ given a sample of trajectories from the approximate transition path process. By importance sampling, we mean the unbiased estimation of observables of the exact process given a sample of paths of the approximate process. Error estimation and importance sampling entail theoretical difficulties that we cannot resolve at this time, but see Section 6 for a simple example where both strategies appear to work well.
4.1 Relative Entropy
In general, if $P$ and $Q$ are probability measures on the same measurable space, one defines the relative entropy (or Kullback–Leibler divergence) of $P$ with respect to $Q$ as follows.
Definition 4.1 (Relative entropy).
If $P$ and $Q$ are probability measures on $(S,\mathscr{S})$ and $P$ is absolutely continuous with respect to $Q$, we define the relative entropy of $P$ with respect to $Q$ by
$D_{\mathrm{KL}}(P\|Q):=\mathds{E}_{Q}\left[\frac{\mathrm{d}P}{\mathrm{d}Q}\log\left(\frac{\mathrm{d}P}{\mathrm{d}Q}\right)\right]=\mathds{E}_{P}\left[\log\left(\frac{\mathrm{d}P}{\mathrm{d}Q}\right)\right].$
If $P$ is not absolutely continuous with respect to $Q$, $D_{\mathrm{KL}}(P\|Q):=\infty$.
The relative entropy is widely used in machine learning and statistics to measure discrepancies between distributions. In particular, maximum likelihood estimation can be understood as relative entropy minimization. The relative entopy is not symmetric and the triangle inequality does not hold, so it is not a metric. However, $D_{\mathrm{KL}}(P\|Q)\geq 0$ for all $P$ and $Q$, and $D_{\mathrm{KL}}(P\|Q)=0$ implies $P=Q$. Moreover, $D_{\mathrm{KL}}(P\|Q)$ relates to some well-known metrics. For example, Pinsker’s inequality bounds the total variation distance in terms of $D_{\mathrm{KL}}$:
$\lVert P-Q\rVert_{\mathrm{TV}}=\max_{A\in\mathscr{S}}\,\lvert P(A)-Q(A)\rvert\leq\sqrt{\frac{1}{2}D_{\mathrm{KL}}(P\|Q)}.$
By abuse of notation, throughout the rest of this work, we will write $\mathds{P}$ and $\mathds{Q}$ for the restrictions of $\mathds{P}$ and $\mathds{Q}$ to the stopping time $\sigma$-algebra $\mathcal{F}_{\tau}$. In other words, we will stop observing the process $Y_{t}$ when it hits $B$ at time $\tau$. By formula (12) in Theorem 3.11, we have
$D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ $=\mathds{E}_{\mathds{P}}\left[\log\left(\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}\right)\right]$
$=\log\zeta-\log\eta$
$\qquad+\int_{\partial A}\log\left(\frac{m(x)}{\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)}\right)\frac{m(x)}{\eta}\,\mathrm{d}S_{A}(x)$
(13) $\qquad+\mathds{E}_{\mathds{P}}\left[\log\tilde{q}(Y_{\tau})-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right]$
and
$D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P})$ $=\mathds{E}_{\mathds{Q}}\left[\log\left(\frac{\mathrm{d}\mathds{Q}}{\mathrm{d}\mathds{P}}\right)\right]$
$=\log\eta-\log\zeta$
$\qquad+\int_{\partial A}\log\left(\frac{\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)}{m(x)}\right)\frac{\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)}{\zeta}\,\mathrm{d}S_{A}(x)$
(14) $\qquad+\mathds{E}_{\mathds{Q}}\left[\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s-\log\tilde{q}(Y_{\tau})\right].$
Both of these relative entropies measure the discrepancy between the exact distribution of transition paths and the distribution with an approximate committor function $\tilde{q}$ in place of $q$.
These formulas simplify under some conditions. First, note that $\log\tilde{q}(Y_{\tau})=0$ when we impose $\tilde{q}=1$ on $\partial B$. Second, note that the integrals over $\partial A$ vanish when
$m(x)=\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon).$
Of course, both conditions hold for the exact committor and reactive flux distribution, and one could impose either in practice when computing approximate committor functions.
Remark 4.2 (Which Entropy?).
In the present work, we focus on $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ instead of $D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P})$, since $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ can be estimated or minimized given a sample of approximate reactive trajectories drawn from $\mathds{P}$. If one instead had a sample of exact reactive trajectories from $\mathds{Q}$, then one would prefer $D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P})$. A sample of exact reactive trajectories could possibly come from a very long molecular dynamics simulation of a not very metastable system or from an enhanced sampling method like parallel replica dynamics [32]. We leave it as an exercise for the reader to show that if $q_{\theta}$ is a smooth parametric family of functions that satisfy Assumption 3 and $\mathds{P}_{\theta}$ is the distribution of the transition path process with $q_{\theta}$ in place of $q$, then $D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P}_{\theta})$ is differentiable under much weaker conditions on $q_{\theta}$ than Assumption 4.3, which implies differentiability of $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$. Moreover, one can easily estimate $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P}_{\theta})$ given a sample from $\mathds{Q}$. Therefore, given a sample from $\mathds{Q}$ and knowledge of the generator $L$, one could train an approximate committor to minimize $D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P}_{\theta})$.
Remark 4.3.
The existence of the change of measure $\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}$ does not imply $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})<\infty$. We note that $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})<\infty$ under the same conditions that imply finite variance of the estimator in Lemma B.2, but with $\mathds{E}_{\mathds{P}}[\tau^{2}]<\infty$ replaced by $\mathds{E}_{\mathds{P}}[\tau]<\infty$. We leave the proof to the reader.
4.2 Selection: Estimating Relative Entropy Differences
Let $\tilde{q}$ and $\bar{q}$ be approximate committor functions for which Assumption 3 holds. Let $\tilde{m}$ and $\bar{m}$ be approximate reactive flux densities supported on $\partial A$ with normalizing constants
[…]
Assume that $\tilde{m}$ and $\bar{m}$ are strictly positive. We define $\tilde{\mathds{P}}$ to be the law of the approximate transition path process corresponding to $\tilde{q}$ and $\tilde{m}$, i.e. the law of a weak solution of (8) with singular drift $2\varepsilon\nabla\log\tilde{q}(Y_{t})$ and initial condition $Y_{0}\sim\tilde{\eta}^{-1}m(x)\,\mathrm{d}S_{A}(x)$. We define $\bar{\mathds{P}}$ similarly, but with $\bar{q}$ and $\bar{m}$ in place of $\tilde{q}$ and $\tilde{m}$. We now explain how to estimate the relative entropy difference
$\delta(\tilde{q},\tilde{m};\bar{q},\bar{m}):=D_{\mathrm{KL}}(\tilde{\mathds{P}}\|\mathds{Q})-D_{\mathrm{KL}}(\bar{\mathds{P}}\|\mathds{Q})$
given samples of paths from $\tilde{\mathds{P}}$ and $\bar{\mathds{P}}$. Using our estimator, one can compare the quality of approximate committor functions. For example, one can monitor the relative entropy difference between initial and successive approximations to the committor to assess progress during training, or one can compare the quality of coarse-grained approximations to the committor depending on different numbers of variables.
Suppose that we have a sample of $\tilde{N}$ paths $\tilde{Y}^{k}_{t}$ drawn from $\tilde{\mathds{P}}$:
$\tilde{Y}^{k}_{t}\overset{\mathrm{i.i.d.}}{\sim}\tilde{\mathds{P}}\text{ for }k=1,\dots,\tilde{N}$.
Under certain conditions on $\tilde{q}$, the sample average
$\tilde{I}_{\tilde{N}}:=\frac{1}{\tilde{N}}\sum_{k=1}^{\tilde{N}}\log\left(\frac{m(\tilde{Y}^{k}_{0})}{\lvert\nabla\tilde{q}(\tilde{Y}^{k}_{0})\rvert\exp(-U(\tilde{Y}^{k}_{0})/\varepsilon)}\right)+\log\tilde{q}(\tilde{Y}^{k}_{\tau})-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(\tilde{Y}^{k}_{s})\,\mathrm{d}s$
is a consistent and unbiased estimator of the term
$\tilde{I}$ $:=\int_{\partial A}\log\left(\frac{m(x)}{\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)}\right)\frac{m(x)}{\eta}\,\mathrm{d}S_{A}(x)$
$\qquad+\mathds{E}_{\mathds{P}_{\theta}}\left[\log\tilde{q}(Y_{\tau})-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right]$
[…]
We have
$\delta(\tilde{q},\tilde{m};\bar{q},\bar{m})=\log\left(\frac{\bar{\eta}}{\tilde{\eta}}\right)+\tilde{I}-\bar{I},$
so it remains to estimate the ratio $\frac{\bar{\eta}}{\tilde{\eta}}$. Analogous problems involving ratios of normalizing constants arise in Bayesian model selection and the calculation of alchemical free energy differences. Refer to [21] for a survey of methods for free energy calculations. We propose to use the Bennett Acceptance Ratio (BAR) method [4] to estimate $\frac{\bar{\eta}}{\tilde{\eta}}$. BAR is in some sense the optimal estimator of $\frac{\bar{\eta}}{\tilde{\eta}}$ given i.i.d. samples from $\tilde{m}\,\mathrm{d}S$ and $\bar{m}\,\mathrm{d}S$. More complex alternatives that involve sampling from multiple distributions, such as the Multistate Bennett Acceptance Ratio (MBAR) method, may be needed in cases where $\tilde{m}$ and $\bar{m}$ differ greatly [28].
Remark 4.4.
In practice, to sample from the approximate reactive flux densities $\tilde{m}$ and $\bar{m}$, one would most likely use a Markov chain Monte Carlo (MCMC) method. Many methods have been developed to sample distributions supported on submanifolds such as $\partial A$. For example, see [21, Sections 3.2.3-4]. Of course, MCMC would produce a correlated sample from the approximate reactive flux distribution, not an independent sample as assumed above.
4.3 Training On the Fly: Minimizing the Relative Entropy while Generating Reactive Trajectories
One could also try to minimize the relative entropy over a family of approximate committor functions by gradient descent. Let $\{q_{\theta};\theta\in\mathbb{R}^{k}\}$ be such a family. Let each approximate committor function correspond to an approximate reactive flux distribution
(15) $m_{\theta}(x):=\eta_{\theta}^{-1}\nabla q_{\theta}(x)\cdot n(x)\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x),$
where
$\eta_{\theta}:=\int_{\partial A}\nabla q_{\theta}(x)\cdot n(x)\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x).$
We define $\mathds{P}_{\theta}$ to be the law of a weak solution of the stochastic differential equation (8) observed up to time $\tau$ for the approximate committor function $q_{\theta}$ and initial condition $Y_{0}\sim m_{\theta}$. In this section, we derive a consistent estimator of $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ given a sample from $\mathds{P}_{\theta}$. We then use the estimator in a stochastic gradient descent method to minimize $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ in Section 6.
For our results in this section, we require much stronger assumptions on $q_{\theta}$ and on the distribution of $\tau$ than those imposed elsewhere; cf. Assumption 4.3. The most significant of these new assumptions is that
(16) $Lq_{\theta}=0\text{ on }\partial A\text{ for all }\theta\in\mathbb{R}^{k}.$
Recall that for the exact committor $q$, or more precisely for the smooth extension guaranteed by Lemma 2.2, we have $Lq=0$ everywhere in $\bar{\mathscr{T}}$, including on $\partial A$; cf. Remark 2.4. In Section 5.1, we explain how to construct a practical families $q_{\theta}$ satisfying (16) when the boundary of $A$ is planar. However, we do not claim that (16) is necessary for differentiability. We simply cannot prove differentiability without imposing such an assumption.
If $Lq_{\theta}=0$ on $\partial A$, then the quantity
$\ell(x;\theta):=\frac{Lq_{\theta}}{q_{\theta}}(x)$
that appears in the change of measure formula (12) extends to a smooth function of $x$ defined on an open neighborhood of $\partial A$, even though we assume $q_{\theta}=0$ on $\partial A$. To see this, let $Lq_{\theta}$ take the place of $\tilde{q}$ in the proof of Lemma 3.1. Similarly,
$\nabla_{\theta}\ell(x;\theta)=\frac{\nabla_{\theta}Lq_{\theta}(x)}{q_{\theta}(x)}-\frac{\nabla_{\theta}q_{\theta}(x)}{q_{\theta}(x)}\frac{Lq_{\theta}(x)}{q_{\theta}(x)}$
also extends to a smooth function on an open neighborhood of $\partial A$, since we have $Lq_{\theta}=q_{\theta}=0$ on $\partial A$ for all $\theta$, hence $\nabla_{\theta}Lq_{\theta}=\nabla_{\theta}q_{\theta}=0$ on $\partial A$. We will assume that both $\ell(x;\theta)$ and $\nabla_{\theta}\ell(x;\theta)$ are not merely smooth on a neighborhood of $\partial A$ but bounded uniformly over all $x$ and $\theta$.
Another new assumption is that, for each $\theta\in\mathbb{R}^{k}$, there is some $\gamma(\theta)>0$ so that
(17) $\mathds{E}_{\mathds{P}_{\theta}}[\exp(\gamma(\theta)\tau)]<\infty;$
that is, the moment generating function of $\tau$ under $\mathds{P}_{\theta}$ is finite on an open interval containing zero. We will not verify (17) for any particular family of approximate committor functions. However, we note that for the exact transition path process in one-dimension, under some rather stringent conditions on the potential energy $U$ and on the sets $A$ and $B$, one can show that $\tau$ converges in distribution as $\varepsilon\rightarrow 0$ to a shifted and scaled Gumbel random variable [7]. Under much more general conditions, escape times from basins of attraction of minima of $U$ are approximately exponentially distributed for the overdamped Langevin dynamics in the limit of small $\varepsilon$ [15]. The moment generating function of a Gumbel or exponential random variable will be finite on an open interval containing zero, but not globally finite, which motivates (17).
We summarize these new assumptions together with a few others below.
{assumption}
We assume that $q_{\theta}(x)$ is infinitely differentiable in both $x$ and $\theta$. We impose the following boundary conditions:
  • $q_{\theta}=0$ for $x\in\partial A$.
  • $q_{\theta}(x)>0$ for $x\in\mathring{\mathscr{T}}\cup\partial B$.
  • $\nabla q_{\theta}(x)\cdot n(x)>0$ for $x\in\partial A$.
  • $Lq_{\theta}(x)=0$ for $x\in\partial A$.
We assume the following uniform bounds:
  • For some $C>0$,
    $\left\lvert\frac{Lq_{\theta}}{q_{\theta}}(x)\right\rvert\leq C\text{ and }\left\lvert\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(x)\right\rvert\leq C$
    for all $x\in\bar{\mathscr{T}}$ and all $\theta\in\mathbb{R}^{k}$.
  • For some $M>0$,
    $M\geq\eta_{\theta}\geq\frac{1}{M}>0\text{ and }\lvert\nabla_{\theta}m_{\theta}\rvert\leq M$
    for all $\theta\in\mathbb{R}^{k}$.
We assume that for each $\theta\in\mathbb{R}^{k}$, there is some $\gamma(\theta)>0$ so that
(18) $\mathds{E}_{\mathds{P}_{\theta}}[\exp(\gamma(\theta)\tau)]<\infty.$
Under Assumption 4.3, the relative entropy is differentiable.
Theorem 4.5.
Let Assumption 4.3 hold, and let $\mathds{P}_{\theta}$ and $m_{\theta}$ be defined as above. The relative entropy $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ is differentiable as a function of $\theta$, and
$\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ $=\operatorname{cov}_{\mathds{P}_{\theta}}\left(\nabla_{\theta}\log q_{\theta}(Y_{\tau})-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(Y_{s})\,\mathrm{d}s,\right$.
$\qquad\qquad\qquad\qquad\left.\log q_{\theta}(Y_{\tau})-\int_{0}^{\tau}\frac{Lq_{\theta}}{q_{\theta}}(Y_{s})\,\mathrm{d}s\right).$
We give the proof in Appendix B. We propose to estimate $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ like any other covariance. Given a sample
$Y^{k}\overset{\mathrm{i.i.d.}}{\sim}\mathds{P}_{\theta}\text{ for }k=1,\dots,N,$
we define sample averages
$J_{N}$ $:=\frac{1}{N}\sum_{k=1}^{N}\log q_{\theta}(Y^{k}_{\tau})-\int_{0}^{\tau}\frac{Lq_{\theta}}{q_{\theta}}(Y^{k}_{s})\,\mathrm{d}s$
$K_{N}$ $:=\frac{1}{N}\sum_{k=1}^{N}\nabla_{\theta}\log q_{\theta}(Y^{k}_{\tau})-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(Y^{k}_{s})\,\mathrm{d}s$
and the sample covariance
(19) $\begin{split}G_{N}&:=\frac{1}{N-1}\sum_{k=1}^{N}\left(\log q_{\theta}(Y^{k}_{\tau})-\int_{0}^{\tau}\frac{Lq_{\theta}}{q_{\theta}}(Y^{k}_{s})\,\mathrm{d}s-J_{N}\right)\\ &\qquad\qquad\qquad\times\left(\nabla_{\theta}\log q_{\theta}(Y^{k}_{\tau})-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(Y^{k}_{s})\,\mathrm{d}s-K_{N}\right).\end{split}$
Assumption 4.3 implies that $G_{N}$ has finite variance. We leave the proof to the reader, noting that it is similar to the proof of Lemma B.2. Using this gradient estimator, one can attempt to minimize $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ by stochastic gradient descent or variants such as Adam [19].
4.4 Importance Sampling and Direct Estimation of Relative Entropy
Using our change of measure formula, one could perhaps use importance sampling to estimate arbitrary expectations over the transition path process $\mathds{Q}$ given a sample of paths from the approximate process $\mathds{P}$. The discussion in this section is speculative. We do not claim that all estimators presented below are robust in practice or even that they would have finite variance under all conditions.
Let $g:\Omega\rightarrow\mathbb{R}$ be $\mathcal{F}_{\tau}$-measurable, i.e. a function of the transition path process observed up to time $\tau$. By (12), assuming for simplicity that $\tilde{q}=1$ on $\partial B$ and $m=\lvert\nabla\tilde{q}\rvert\exp(-U/\varepsilon)$, we have
$\mathds{E}_{\mathds{Q}}[g]$ $=\mathds{E}_{\mathds{P}}\left[Z_{\tau}^{-1}g\right]$
$=\frac{\eta}{\zeta}\mathds{E}_{\mathds{P}}\left[\exp\left(\int_{0}^{\tau(\omega)}\frac{L\tilde{q}}{\tilde{q}}(Y_{s}(\omega))\,\mathrm{d}s\right)g(\omega)\right],$
which suggests the self-normalized importance sampling estimator
(20) $\bar{g}_{N}=\frac{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)g(Y^{k})}{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)},$
of $\mathds{E}_{\mathds{Q}}[g]$, where $Y^{k}\overset{\mathrm{i.i.d.}}{\sim}\mathds{P}\text{ for }k=1,\dots,N$. Experience with other estimators related to Girsanov’s theorem suggests that such a naïve estimator will not be robust. Some successes have been attained in the importance sampling of paths of diffusion processes by roughly similar means [34, 11, 30]. We hope that our results will facilitate the development of analogous methods for the transition path process. However, we note that [34, 11, 30] treat only paths observed up to a fixed finite time, not up to an unbounded stopping time as in our work. This is a significant challenge. In fact, since the stopping time is expected to have a moment generating function defined only a finite interval, it is not clear to us that the variance of $\bar{g}_{N}$ must be finite even if one assumes that $L\tilde{q}/\tilde{q}$ is bounded. By contrast, note that our estimators of relative entropy differences and gradients do have finite variance under plausible assumptions. Nonetheless, in Section 6, we present numerical experiments related to a simple problem for which the self-normalized importance sampling estimator produces good results.
One might also consider the possibility of estimating the relative entropy $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$, not a relative entropy difference, for a given approximate committor function. The crux of the problem is to estimate the ratio of normalizing constants
$\frac{\zeta}{\eta}$
given a sample of transition paths from $\mathds{P}$. In Section 4.2, we explain how to estimate all other terms comprising the relative entropy. As above, we have
$\frac{\zeta}{\eta}$ $=\mathds{E}_{\mathds{P}}\left[\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)\right],$
which suggests the estimator
(21) $\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)$
for $\zeta/\eta$, where $Y^{k}\overset{\mathrm{i.i.d.}}{\sim}\mathds{P}\text{ for }k=1,\dots,N$. We do not expect this estimator to be any more robust than the more general importance sampling estimator above.
5 Numerical Methods
In this section, we address some obstacles to the practical implementation of the estimators proposed above. In Section 5.1, we explain how to construct families of committor functions that satisfy Assumption 3 and the boundary condition $L\tilde{q}=0$ on $\partial A$ that arises in Section 4.3. In Section 5.2, we explain how to handle the possibly singular integrand $\frac{L\tilde{q}}{\tilde{q}}$ that appears in our change of measure formula (12).
5.1 Representing the Committor Function
We devise a practical means of representing approximate committor functions that satisfy Assumption 3 and the condition
$L\tilde{q}=0\text{ on }\partial A$
appearing in Assumption 4.3. In our approach, one must first choose a computable function $T:\mathscr{T}\rightarrow\mathbb{R}$ with the following properties:
  • $T$ extends to a smooth function on an open neighborhood of $\overline{\mathscr{T}}$.
  • $\nabla T(x)\cdot n(x)>0$ for all $x\in\partial A$.
  • $T(x)=0$ for all $x\in\partial A$.
  • $T(x)>0$ for all $x\in\mathscr{T}\setminus\partial A$.
For example, if $A=\{x\in\mathbb{R}^{d};\xi(x)\leq a\}$ were a sublevel set of some function $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}$, one could try $T=\xi-a$. Observe that all of the above conditions would hold at least locally in a neighborhood of $\partial A$ if $\lvert\nabla\xi(x)\rvert$ were positive on $\partial A$.
In our numerical experiments, the approximate committor takes the form
$\tilde{q}=T\exp(w),$
where $T$ is as above and $w:\mathscr{T}\rightarrow\mathbb{R}$ belongs to some convenient family of smooth functions. Any $\tilde{q}$ of this form satisfies Assumption 3. In particular, the normal derivative of $\tilde{q}$ is positive over $\partial A$, since
$\nabla\tilde{q}(x)\cdot n(x)$ $=\big\{\nabla T(x)\exp(w(x))+\nabla w(x)T(x)\exp(w(x))\big\}\cdot n(x)$
$=\nabla T(x)\cdot n(x)\exp(w(x))$
$>0$
for $x\in\partial A$. Moreover, for any $T$ meeting the conditions above, the exact committor $q$ can be written in the form $q=T\exp(w)$ for $w=\log\left(q/T\right)$, and this $w$ extends to a smooth function on an open neighborhood of $\Omega$ by Lemma 3.1.
The condition $L\tilde{q}=0$ on $\partial A$ appearing in Assumption 4.3 is equivalent with a Neumann boundary condition on $w$. We have
(22) $L\tilde{q}=\left\{LT+2\varepsilon\nabla w\cdot\nabla T+T\left(Lw+\varepsilon\lvert\nabla w\rvert^{2}\right)\right\}\exp(w).$
Since $T=0$ and $\nabla T\cdot n>0$ on $\partial A$, $\nabla T=\lvert\nabla T\rvert n$ on $\partial A$, and therefore (22) implies that $L\tilde{q}=0$ on $\partial A$ if and only if $w$ satisfies the Neumann boundary condition
(23) $\nabla w(x)\cdot n(x)=-\frac{1}{2\varepsilon}\frac{LT(x)}{\lvert\nabla T(x)\rvert}$
for $x\in\partial A$. Of course, the difficulty of imposing such a boundary condition in practice depends on the geometry of $\partial A$.
5.2 Alternative Expression for the Improper Integral
Observe that $\frac{L\tilde{q}}{\tilde{q}}$ must be singular on $\partial A$ unless $L\tilde{q}=0$ on $\partial A$. Moreover, to compute $L\tilde{q}$ requires computing $\Delta\tilde{q}$. If $\tilde{q}$ is an artifical neural network and the state space is high-dimensional, then it may be inefficient to compute $\Delta\tilde{q}$ by automatic differentiation. See [16] for an explanation of the problem and for a general strategy that can be used to efficiently estimate $\Delta\tilde{q}$ in a stochastic gradient descent method at the cost of increasing the variance of the gradient estimates. Here, we derive a convenient alternative expression for the improper integral term
$\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s$
in the relative entropy (13). We eliminate the computationally undesirable dependence on $\Delta\tilde{q}$. We also explain why the integral is finite even though the integrand is singular. Lemma 5.1 below is the crucial result.
Lemma 5.1.
Let $S:\mathscr{T}\rightarrow\mathbb{R}$ be a function for which Assumption 3 holds. Let $w=\log\frac{\tilde{q}}{S}$, so
$\tilde{q}=S\exp(w).$
We have
$\int_{0}^{t}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s=w(Y_{t})-w(Y_{0})$ $+\int_{0}^{t}\frac{LS}{S}(Y_{s})-\epsilon|\nabla w(Y_{s})|^{2}\,\mathrm{d}s$
$-\int_{0}^{t}\sqrt{2\epsilon}\nabla w(Y_{s})\cdot\mathrm{d}W_{s},$
where $W_{t}$ is the $\mathds{P}$-Brownian motion in Theorem 3.11.
We give the proof in Appendix B. For theoretical purposes, we may take $S=q$ in Lemma 5.1, so that $w=\log r=\log(\tilde{q}/q)$, $LS=0$ everywhere in $\bar{\mathscr{T}}$ including on $\partial A$, and the $\frac{LS}{S}$ term vanishes. In that case, one recovers a formula for the change of measure that is similar to (10). We use this formula in the proofs of some results in Section 4. For computational purposes, one can choose $S$ so that $\Delta S$ is easy to compute explicitly without resorting to automatic differentiation, and then one need only compute $\nabla\tilde{q}$, not $\Delta\tilde{q}$, when evaluating $\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{t})\,\mathrm{d}t$. Note that if $w$ is a neural network, then $\Delta\tilde{q}$ will most likely be intractable, but $\nabla\tilde{q}$ can be computed efficiently by backpropagation.
We also note that the formula implies that the integral $\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{t})\,\mathrm{d}t$ is finite with probability one even though the integrand is singular on $\partial A$. In fact, assuming that $r$ and $\nabla\log r$ are bounded,
$\mathds{E}_{\mathds{P}}\left[\int_{0}^{h}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right]=O(\sqrt{h})\text{ and }\operatorname{var}_{\mathds{P}}\left(\int_{0}^{h}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)=O(\sqrt{h})$
in the limit as $h\rightarrow 0$. This suggests that it should be feasible to devise an accurate numerical integrator for the calculation of $\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s$.
Remark 5.2.
We warn the reader that since $W_{t}$ depends on $\tilde{q}$, one has to be rather careful in attempts to use this formula to simplify the calculation of derivatives of the relative entropy.
6 Evaluating a One-Dimensional Approximation of a Toy Two-Dimensional System
Here, we illustrate how the methods devised above might be used to assess the quality of a low-dimensional approximation to a committor function, to train an approximate committor function, and to compute statistics of reactive trajectories. For readability, we do not describe all aspects of the implementation of our numerical methods in this section; see Appendix C for details.
6.1 Toy Two-Dimensional System
We consider a two-dimensional system, where
$A=\{x\in\mathbb{R}^{2};x_{1}\leq a\}\text{ and }B=\{x\in\mathbb{R}^{2};x_{1}\geq b\}$
for
$a=-0.75\text{ and }b=0.85.$
We define $U_{1}:\mathbb{R}\rightarrow\mathbb{R}$ by
$U_{1}(x)=3\left(x^{2}+\frac{1}{20}\right)\left(5(x^{2}-1)^{2}+\frac{x}{2}\right)$
and $U:\mathbb{R}^{2}\rightarrow\mathbb{R}$ by
$U(x_{1},x_{2})=U_{1}(x_{1})+x_{2}^{2}+x_{2}(x_{1}-0.515)^{2}.$
We take
$\varepsilon=\frac{10}{13}.$
A contour plot of $U$ with the reagent and product sets indicated appears in Figure 2.
6.2 One-Dimensional Approximation to the Committor
Observe that only the single term $x_{2}(x_{1}-0.515)^{2}$ in $U$ couples $x_{1}$ with $x_{2}$. If one eliminates that term, then the overdamped Langevin dynamics decouples into independent equations for the coordinate components. The first component satisfies
$\mathrm{d}X_{t,1}=-U^{\prime}_{1}(X_{t,1})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t,1}.$
In that case, the committor $q_{1}$ is a function of $x_{1}$ alone, and it solves
$-U^{\prime}_{1}(x_{1})q_{1}^{\prime}(x_{1})+\epsilon q_{1}^{\prime\prime}(x_{1})$ $=0,$
$q_{1}(a)$ $=0,$
$q_{1}(b)$ $=1.$
Using integrating factors yields
(24) $q_{1}(x_{1})=\frac{\int_{a}^{x_{1}}\exp\left({\frac{U_{1}(x)}{\epsilon}}\right)\,\mathrm{d}x}{\int_{a}^{b}\exp\left({\frac{U_{1}(x)}{\epsilon}}\right)\,\mathrm{d}x}.$
We calculated $q_{1}$ by quadrature using the integrate module in SciPy [31]; cf. Figure 3. In the numerical experiments below, we assess the quality of $q_{1}$ as an approximation of the exact committor $q$ and we train an improved approximation.
Figure 3: The coarse-grained approximate committor $q_{1}$ and the corresponding effective potential $U-2\varepsilon\log q_{1}$. In the left pane, we plot $q_{1}$. In the middle, we plot the effective potential $U-2\varepsilon\log q_{1}$ corresponding to $q_{1}$. The gradient of the effective potential is the biasing force on reactive trajectories generated using $q_{1}$. For comparison, in the third panel, we plot the effective potential $U-2\varepsilon\log q_{\rm{fem}}$, where $q_{\rm{fem}}$ is a very accurate finite element approximation to the committor.
For comparison with $q_{1}$ and to validate our results, we computed a very accurate approximation of the committor by a finite element method, minimizing the Ritz functional
(25) $R(\tilde{q}):=\int_{[a,b]\times[-4,4]}\lvert\nabla\tilde{q}(x)\rvert^{2}\exp(-U(x)/\varepsilon)\,\mathrm{d}x$
on the domain $a,b]\times[4,4]$ with Dirichlet boundary conditions $q=0$ on $\partial A$ and $q=1$ on $\partial B$ and the natural zero Neumann boundary condition on the top and bottom of the domain where $x_{2}=\pm 4$. The Neumann condition corresponds to confining the overdamped Langevin dynamics to the domain by reflection. The results appear in the middle panel of Figure 4. Observe that the exact committor depends significantly on $x_{2}$.
6.3 A Numerical Integrator Based on Operator Splitting
Estimating the relative entropy and its gradient requires a numerical integrator for the approximate transition path process. To devise such an integrator, we write the approximate committor function in the form
$\tilde{q}(x)=T(x)\exp(w(x)),$
where
$T(x)=\frac{x-a}{b-a}.$
By Lemma 3.1, any $\tilde{q}$ that satisfies Assumption 3 can be written in the above form for some smooth function $w$; cf. the more general discussion in Section 5.1. For such $\tilde{q}$, the approximate transition path equation is
$\mathrm{d}Y_{t}$ $=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log\tilde{q}(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
$=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla w(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log T(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
(26) $=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla w(Y_{t})\,\mathrm{d}t+\frac{2\varepsilon}{Y_{t,1}-a}e_{1}\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}.$
Here, $Y_{t,1}$ denotes the first coordinate component of the two-dimensional vector $Y_{t}$, and $e_{1}=(1,0)\in\mathbb{R}^{2}$ is the first standard basis vector. We propose an integrator based on a formal splitting of the generator of $Y_{t}$.
Consider the equation
(27) $\mathrm{d}X_{t}=\frac{2\varepsilon}{X_{t,1}-a}e_{1}\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
corresponding to the last two terms in (26). This will be the first piece in our splitting. Equation (27) can be simulated exactly: The first coordinate component is a three-dimensional Bessel process and the second is a Brownian motion. To be precise, if $W_{t}$ is a three-dimensional Brownian motion, then $R_{t}=\lvert W_{t}\rvert$ is said to be a three-dimensional Bessel process, and
(28) $\mathrm{d}R_{t}=\frac{1}{R_{t}}\mathrm{d}t+\mathrm{d}\hat{W}_{t}$
for some one-dimensional Brownian motion $\hat{W}_{t}$ [26]. Equation (28) holds even when $R_{0}=0$. Note the similarity between the drift terms in (28) and (27). To construct a weak solution of (27), we may take
(29) $X_{t}=(R_{2\varepsilon t}+a,W_{2\varepsilon t}),$
where $W_{t}$ is a one-dimensional Brownian motion that is independent of $R_{t}$. We observe that $X_{t}$ remains outside $A$ for all time, since the Bessel process $R_{t}$ must be nonnegative and is in fact strictly positive for $t>0$.
Increments of the Bessel-like process (27) can be simulated conveniently. To see this, let $X_{t}$ solve (27) with $X_{0}=x$ where $x_{1}\geq a$. Because of the relationship between $X_{t}$ and the three-dimensional Bessel process, $X_{\Delta t,1}$ has the same law as
$\lVert(x_{1}-a,0,0)+\sqrt{2\varepsilon\Delta t}\xi\rVert+a$
where $\xi\sim N(0,I)$ is a vector of three independent standard normal random variables. We note that, in this formula, instead of $(x_{1}-a,0,0)$, one could choose any vector of length $x_{1}-a$, since increments of three-dimensional Brownian motion have a distribution that is symmetric under rotation.
Our integrator alternates increments of the Bessel-like process (27) with increments of the ordinary differential equation
(30) $\mathrm{d}Y_{t}=(-\nabla U(Y_{t})+2\varepsilon\nabla w(Y_{t}))\,\mathrm{d}t$
computed by the Euler method. See Algorithm 1 for details. We do not prove convergence of the integrator in this work, but see Sections 6.5 and 6.6 for numerical studies that suggest convergence.
Algorithm 1 An integrator based on operator splitting.
  1. 1.
    Fix a time step $\Delta t>0$.
  2. 2.
    Suppose that $\tilde{Y}_{n}$ has been calculated as an approximation of $Y_{n\Delta t}$. Generate a vector
    $\xi_{n}=(\xi_{n,1},\xi_{n,2},\xi_{n,3})\sim\mathrm{N}(0,I)$
    of three independent standard normal random variables. Let $\xi_{n}$ be independent of all previously generated random variables. Set
    $\tilde{Y}_{n+\frac{1}{2},1}=\lVert(\tilde{Y}_{n,1}-a,0,0)+\sqrt{2\varepsilon\Delta t}\xi_{n}\rVert+a.$
    Let $\eta_{n}\sim\mathrm{N}(0,1)$ be a standard normal random variable that is independent of all previously generated random variables, including $\xi_{n}$. Set
    $\tilde{Y}_{n+\frac{1}{2},2}=\tilde{Y}_{n,2}+\sqrt{2\varepsilon\Delta t}\eta_{n}.$
    Here, $\tilde{Y}_{n+\frac{1}{2}}$ has the same law as the Bessel-like process (27) at time $t=\Delta t$ with the initial condition $X_{0}=\tilde{Y}_{n}$.
  3. 3.
    Set $\tilde{Y}_{n+1}=\nabla w(\tilde{Y}_{n+\frac{1}{2}})\Delta t+\tilde{Y}_{n+\frac{1}{2}}$. That is, perform one step of the Euler method for the ordinary differential equation (30). Return to step (2).
6.4 Training an Improved Approximate Committor
Here, we embed the low-dimensional or coarse-grained approximation $q_{1}$ in a parametric family of approximate committor functions and minimize the relative entropy by the stochastic gradient descent method of Section 4.3. We obtain both an improved approximation to the committor and also a measure of the error of the coarse-grained approximation $q_{1}$. To be precise, we let
(31) $q_{\theta}(x)=q_{1}(x)\exp(w_{\theta}(x)(b-x_{1})),$
where $w_{\theta}$ is a tensor product of cubic B-splines on a uniform $4\times 16$ grid covering $a,b]\times[-3,3]$. Every member of this family satisfies Assumption 3 and has $q_{\theta}=1$ on $\partial B$.
We minimized $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ over $\theta$, which amounts to fitting the difference $\log q-\log q_{1}$ by a spline using the relative entropy as a loss function. We used the Adam optimizer [19] and our gradient estimator (19). Only two points related to the implementation are relevant to the discussion below: We approximated the integrals in (19) by right-handed Riemann sums over the discrete trajectories, so the first point at which $\frac{Lq_{\theta}}{q_{\theta}}$ is evaluated is $X_{\Delta t}$. We took $\Delta t=0.005$, which appears to be very close to the maximum for which the integrator is stable. See Appendix C for additional details.
Figures 4 and 5 summarize the results. Figure 4 shows that the approximate committor $q_{\rm{sgd}}$ computed by gradient descent is at least qualitatively accurate. The large errors observed near the top and bottom of the domain occur in regions of extremely small probability under the Boltzmann distribution and might be considered irrelevant for most purposes. In any case, we cannot expect an accurate estimate of the committor in those regions when our reactive trajectories visit them so rarely. However, since we use only a small $4\times 16$ grid in our spline approximation, relative errors approach $10\%$ even in some regions that reactive trajectories visit frequently. Note that relative errors, or equivalently errors in logarithms, are most relevant here, since the singular drift term $2\varepsilon\nabla\log\tilde{q}$ is a function of $\log\tilde{q}$.
On the other hand, Figure 5 suggests that if one desires primarily the singular drift for simulating reactive trajectories, then $q_{\rm{sgd}}$ should be considered quantitately accurate. The relative entropy $D_{\mathrm{KL}}(\mathds{P}_{q_{\rm{sgd}}}\|\mathds{Q})$, where $\mathds{P}_{q_{\rm{sgd}}}$ is the law of the transition path process with $q_{\rm{sgd}}$ in place of $q$, is approximately $0.011$; cf. the right panel in Figure 5. Therefore, in light of results such as Pinsker’s inequality, $q_{\rm{sgd}}$ should produce reactive trajectories having nearly the correct distribution. We present a study of crossover times below to test whether this is in fact the case. We note that the Ritz form $R(q_{\rm{sgd}})$ is also very close to the minimum value. For the committor $q_{\rm{fem}}$ computed by the finite element method, $R(q_{\rm{fem}})=0.06612$, and $R(q_{\rm{sgd}})$ is roughly $0.06624$, cf. the right panel in Figure 4.
Figure 4: Approximate committor function calculated by fitting the difference between $\log q$ and $\log q_{1}$ by a spline. The left pane shows the approximate committor function. The middle shows a very nearly exact committor calculated by the finite element method. The right shows the percent relative error.
Figure 5: Convergence of a stochastic gradient descent method for fitting the difference between $\log q$ and $\log q_{1}$ by a spline. The blue curve in the left pane is an estimate of $D_{\mathrm{KL}}(\mathds{P}_{\theta_{n}}|\mathds{Q})$ computed using (21) with the average taken over the batch of $64$ trajectories generated in each step of gradient descent. The blue curve in the right pane is the same, but smoothed by averaging over the preceding $32$ gradient descent steps. The red curves are quadrature estimates of the Ritz form $R(q_{\theta_{n}})$ defined in (25). The left pane shows the first $512$ out of $1024$ steps of gradient descent, and the right shows the last $512$. The blue scale on the vertical axis relates to the entropy estimates, and the red scale to the Ritz forms.
6.5 Crossover Times, Numerical Error, and Importance Sampling
Here, we study crossover times of numerical reactive trajectories. Our primary objective is to validate the operator splitting integrator. We also demonstrate that the importance sampling formula (20) can be used to correct for errors in approximate committors to produce unbiased estimates of statistics of the exact distribution of transition paths. We calculated the very precise estimate
$\mathds{E}_{\mathds{Q}}[\tau]_{\rm{fem}}=1.153$
of the expected crossover time by substituting $q_{\rm{fem}}$ for $q$ in the formula
(32) $\mathds{E}_{\mathds{Q}}[\tau]=\frac{1}{\varepsilon R(q)}\int_{\mathscr{T}}q(x)(1-q(x))\exp(-U(x)/\varepsilon)\,\mathrm{d}x,$
which appears in Proposition 1.8 in [24], estimating the integrals by quadrature. Here, $R(q)$ is the Ritz form (25). We then generated samples of $N=2^{15}$ approximate reactive trajectories for various values of $\Delta t$ and for the two approximate committor functions $q_{1}$ and $q_{\rm{sgd}}$. We computed average crossover times for each sample both with and without importance sampling. The results appear in Table 1. With importance sampling, for either $q_{1}$ or $q_{\rm{sgd}}$, our integrator produces estimates of the expected crossover time that are very close to $\mathds{E}_{\mathds{Q}}[\tau]_{\rm{fem}}$, especially for $\Delta t\leq 0.001$. Without importance sampling, both $q_{1}$ and $q_{\rm{sgd}}$ produce crossover times that are too long on average, but note that the crossover times for $q_{\rm{sgd}}$ are much closer to $\mathds{E}_{\mathds{Q}}[\tau]_{\rm{fem}}$, indicating that training by stochastic gradient descent did result in significantly better approximate reactive trajectories, as one would expect. For the pooled sample with $\Delta t\leq 0.001$, the error in the expected crossover time for $q_{\rm{sgd}}$ is no more than $1\%$, whereas the error for $q_{1}$ is around $23\%$.
Table 1: Crossover times estimated by numerical integration of the transition path process with an approximate committor. All figures appear in the form $\bar{\tau}\pm s$, where $\bar{\tau}$ is the average crossover time for a sample of $N=2^{15}$ approximate reactive trajectories and $s$ is an estimate of the standard deviation of $\bar{\tau}$. For importance sampling estimates, the standard deviation was computed using the delta method; cf. Appendix C. Figures appear in boldface when the exact value $\mathds{E}_{\mathds{Q}}[\tau]_{\rm{fem}}=1.153$ computed using the finite element method lies within one standard deviation of $\bar{\tau}$. The bottom row contains averages of pooled samples for the smallest three values of $\Delta t$. Note that the pooled averages computed with importance sampling are within roughly $0.5\%$ of the exact value. Here, the sample averages do not depend strongly on $\Delta t$, but compare Table 2.
$\Delta t$ $q_{1}$ without IS $q_{1}$ with IS $q_{\rm{sgd}}$ without IS $q_{\rm{sgd}}$ with IS
5.0e-03 1.438 $\pm$ 0.007 1.166 $\pm$ 0.008 1.182 $\pm$ 0.006 1.170 $\pm$ 0.006
1.0e-03 1.406 $\pm$ 0.007 1.145 $\pm$ 0.008 1.159 $\pm$ 0.005 1.150 $\pm$ 0.006
2.0e-04 1.417 $\pm$ 0.007 1.150 $\pm$ 0.008 1.167 $\pm$ 0.006 1.159 $\pm$ 0.006
5.0e-05 1.415 $\pm$ 0.007 1.152 $\pm$ 0.008 1.157 $\pm$ 0.005 1.148 $\pm$ 0.005
total 1.413 $\pm$ 0.004 1.149 $\pm$ 0.005 1.161 $\pm$ 0.003 1.153 $\pm$ 0.003
6.6 Estimates of the Entropy and Entropy Differences
Finally, we demonstrate the most obvious way to use our results to evaluate the quality of an approximate committor function: We estimate $D_{\mathrm{KL}}(\mathds{P}_{q_{1}}\|\mathds{Q})$ by the method of Section 4.4. The results appear in Table 2. We used the same samples of reactive trajectories as in our study of crossover times. Observe that when estimating entropies instead of crossover times, the numerical error for large values of $\Delta t$ appears to be more significant. We suppose that this is to be expected given our naïve strategy of estimating the singular integral $\int_{0}^{\tau}\frac{Lq_{1}}{q_{1}}(Y_{s})\,\mathrm{d}s$ by a right-handed Riemann sum.
To validate our numerical estimates of the singular integral and related quantities such as $D_{\mathrm{KL}}(\mathds{P}_{q_{1}}\|\mathds{Q})$, we estimated the ratio of normalizing constants $\frac{\zeta}{\eta}$ by (21), we calculated $\eta=\int_{\partial A}\lvert\nabla q_{1}\rvert\exp(-U/\varepsilon)\mathrm{d}S$ more-or-less exactly by quadrature, and then we took the product to estimate $\zeta$; cf. the last column of Table 2. For comparison, we calculated $\zeta$ by the formula $\zeta=R(q)$ from Proposition 1.8 in [24] with $q_{\rm{fem}}$ in place of $q$, which yields $\zeta_{\rm{fem}}=0.06612$. Roughly speaking, errors decrease from $10\%$ to $1\%$ as the time step $\Delta t$ decreases by a factor of $100$, which is consistent with weak convergence of order one half.
We note that numerical errors in the estimation of singular integrals do not seem to strongly affect the convergence of stochastic gradient descent. Recall that we used $\Delta t=0.005$ when training $q_{\rm{sgd}}$, which is the maximum in Table 2 and close to the maximum for which the integrator is stable. However, the value of the Ritz form for $q_{\rm{sgd}}$ was very nearly optimal after training: $R(q_{\rm{sgd}})=0.06624$ with the finite element method producing only $R(q_{\rm{fem}})=0.06612$, and the estimated relative entropy also decreased quite quickly during training. We tried training with smaller values of $\Delta t$, obtaining results that were only comparable, not improved, when assessed either in terms of relative entropies or Ritz forms.
As explained in Section 4.4, we do not expect the estimator (21) of $\zeta/\eta$ to be completely reliable in general since it is based on exponential averages. Therefore, for more difficult problems, if one has only a poor approximation to the committor, one might prefer to estimate entropy differences as in Section 4.2, since to compute differences does not require any exponential averages. For example, one could estimate $D_{\mathrm{KL}}(\mathds{P}_{q_{1}}\|\mathds{Q})-D_{\mathrm{KL}}(\mathds{P}_{q_{\rm{sgd}}}\|\mathds{Q})$ to measure how much training has improved on the initial estimate $q_{1}$ of the committor. For our simple toy problem, the results would be redundant with Table 2, since (21) works quite well even for the relatively poor approximate committor $q_{1}$.
Table 2: Estimates of the entropy, singular integral, and normalizing constant based on samples of reactive trajectories biased by $q_{1}$. Here, $\bar{\zeta}$ is an estimator of the normalizing constant $\zeta=R(q)$ based on sample averages of singular integrals. For comparison, $R(q_{\rm{fem}})=0.06612$.
$\Delta t$ $D_{\mathrm{KL}}(\mathds{P}_{q_{1}}\|\mathds{Q})$ $\mathds{E}_{\mathds{P}_{q_{1}}}\left[\int_{0}^{\tau}\frac{Lq_{1}}{q_{1}}(Y_{s})\,\mathrm{d}s\right]$ $\bar{\zeta}$
5.0e-03 0.5970 $\pm$ 0.0066 -1.2934 $\pm$ 0.0083 0.0724 $\pm$ 0.0004
1.0e-03 0.6243 $\pm$ 0.0070 -1.3716 $\pm$ 0.0085 0.0689 $\pm$ 0.0004
2.0e-04 0.6439 $\pm$ 0.0071 -1.4225 $\pm$ 0.0087 0.0667 $\pm$ 0.0003
5.0e-05 0.6486 $\pm$ 0.0074 -1.4309 $\pm$ 0.0086 0.0665 $\pm$ 0.0004
7 Conclusion
We derive a change of measure formula that establishes well-posedness of the transition path equation with an approximate committor $\tilde{q}$ in place of the exact committor $q$. Based on the change of measure, we propose a stochastic gradient descent method that minimizes the relative entropy loss $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ to train an approximate committor on the fly while computing reactive trajectories. We are interested in methods of this type for several reasons: First, techniques like adaptive multilevel splitting and PINNs require the user to specify either a reaction coordinate or collocation points in advance, which may be difficult in practice. We compute the committor, which is in many respects the ideal reaction coordinate, and instead of collocation points we generate reactive trajectories. We do require some initial approximation to the committor. We suggest that one might choose the initial approximation so that in the early stages of training the reactive trajectories amount to steered molecular dynamics trajectories with a singular, time-independent biasing force pushing on a few atoms to induce transitions to occur rapidly. Second, our loss function $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ serves as a convenient and interpretable measure of the error in the distribution of reactive trajectories. It seems to us that it can be very difficult to assess the quality of solutions to high-dimensional partial differential equations obtained by PINNs and similar methods, since the usual a posteriori error estimates for low-dimensional systems do not apply. We must be careful not to claim too much here, since estimating the normalizing constant ratio involved in $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ may not always be easy, so one might prefer to work with relative entropy differences instead of relative entropies, cf. Section 4.4. However, we obtained good results for the toy problem in Section 6.6. Third, in practice, neither the committor nor a sample of reactive trajectories suffice in themselves to completely characterize rare transitions in complex systems, and one might wish to compute both at once.
In this article, our objectives were to provide a theoretical foundation for approximations to the transition path process and to demonstrate the feasibility of computational methods for simultaneously training the committor and computing reactive trajectories. We obtained good numerical results for a two-dimensional toy problem, but to apply our proposed methods to complex problems of practical interest would entail a significant amount of additional work, including the development of a numerical integrator for the case where $A$ has a curved boundary. We also note that while the numerical studies in Sections 6.5 and 6.6 suggest that the integrator is convergent, it appears that the estimates of the singular integrals arising in our formulas for the relative entropy and change of measure are only of weak order one half, so a higher order method might be desirable. Finally, it seems to us that a more detailed comparison of the relative entropy with other losses should be pursued, and so should a comparison of our method with other methods for the calculation of approximate committor functions.
Appendix A Proofs of Results in Section 3
Proof A.1 (Proof of Theorem 3.11).
Since we assume the Novikov condition, by Lemma 3.8,
$M_{t}=\frac{\lvert\nabla q(Y_{0})\rvert}{\lvert\nabla\tilde{q}(Y_{0})\rvert}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}\exp\left(-\int_{0}^{t}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)$
is an $\mathcal{F}_{t}$-martingale under $\mathds{Q}$ with $\mathds{E}_{\mathds{Q}}[M_{t}]=1$ for $t\geq 0$. It follows that
$Z_{t}$ $=\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{t}$
is a martingale with $\mathds{E}_{\mathds{Q}}[Z_{t}]=1$: For $0\leq s\leq t$, we have
$\mathds{E}_{\mathds{Q}}[Z_{t}|\mathcal{F}_{s}]$ $=\mathds{E}_{\mathds{Q}}\left[\left.\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{t}\right|\mathcal{F}_{s}\right]$
$=\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\mathds{E}_{\mathds{Q}}[M_{t}|\mathcal{F}_{s}]$
$=\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{s}.$
$=Z_{s},$
Moreover,
$\mathds{E}_{\mathds{Q}}[Z_{t}]$ $=\mathds{E}_{\mathds{Q}}\left[\mathds{E}_{\mathds{Q}}\left[\left.\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{t}\right|\mathcal{F}_{0}\right]\right]$
$=\mathds{E}_{\mathds{Q}}\left[\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\mathds{E}_{\mathds{Q}}[M_{t}|\mathcal{F}_{0}]\right]$
$=\mathds{E}_{\mathds{Q}}\left[\frac{\eta^{-1}m(Y_{0})}{\zeta^{-1}\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}M_{0}\right]$
$=1,$
since $M_{0}=1$ and $Y_{0}\sim\zeta^{-1}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)$ under $\mathds{Q}$.
We now show that $Z_{\tau}$ is the density of $\mathds{P}$ with respect to $\mathds{Q}$ restricted to the stopping time $\sigma$-algebra $\mathcal{F}_{\tau}$. Since $Z_{t}$ is a martingale, the optional stopping theorem [26, Chapter 2, Theorem 3.2] implies that for any fixed $t>0$,
$Z_{t\wedge\tau}=\mathds{E}_{\mathds{Q}}[Z_{t}|\mathcal{F}_{t\wedge\tau}].$
(Here, $t\wedge\tau$ denotes the minimium of $t$ and $\tau$.) Let $A\in\mathcal{F}_{\tau}$. Since $\mathds{P}[\tau<\infty]=1$, we have
$\mathds{P}[A]=\lim_{N\rightarrow\infty}\mathds{P}[A\cap\{\tau\leq N\}].$
Now each of the sets
$A_{N}:=A\cap\{\tau<N\}$
is in $\mathcal{F}_{\tau\wedge N}\subset\mathcal{F}_{N}$, since $A\in\mathcal{F}_{\tau}$. (Here, $\tau\wedge N$ means the minimum of $\tau$ and $N$.) Therefore,
$\mathds{P}[A]$ $=\lim_{N\rightarrow\infty}\mathds{P}[A_{N}]$
(since $\mathds{P}[\tau<\infty]=1$)
$=\lim_{N\rightarrow\infty}\int_{A_{N}}\mathds{P}(\mathrm{d}\omega)$
$=\lim_{N\rightarrow\infty}\int_{A_{N}}Z_{N}(\omega)\mathds{Q}(\mathrm{d}\omega)$
(since $A_{N}\in\mathcal{F}_{N}$ and $Z_{N}=\left.\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}\right\rvert_{\mathcal{F}_{N}}$)
$=\lim_{N\rightarrow\infty}\int_{A_{N}}\mathds{E}_{\mathds{Q}}[Z_{N}(\omega)|\mathcal{F}_{\tau(\omega)\wedge N}]\mathds{Q}(\mathrm{d}\omega)$
(since $A_{N}\in\mathcal{F}_{\tau\wedge N}$)
$=\lim_{N\rightarrow\infty}\int_{A_{N}}Z_{\tau(\omega)\wedge N}(\omega)\mathds{Q}(\mathrm{d}\omega)$
(by optional stopping, as discussed above)
$=\lim_{N\rightarrow\infty}\int_{A_{N}}Z_{\tau(\omega)}(\omega)\mathds{Q}(\mathrm{d}\omega)$
(since for $\omega\in A_{N}$, $\tau(\omega)\wedge N=\tau(\omega)$)
$=\int_{A}Z_{\tau(\omega)}(\omega)\mathds{Q}(\mathrm{d}\omega),$
(since $\mathds{Q}[\tau<\infty]=1$)
and so $Z_{\tau}$ is the density of $\mathds{P}$ with respect to $\mathds{Q}$ on $\mathcal{F}_{\tau}$.
Appendix B Proofs of Results in Section 4
First, we prove Lemma 5.1 which provides an alternative expression for the improper integral in our change of measure formula.
Proof B.1 (Proof of Lemma 5.1).
By Lemma 3.1, $w$ is smooth. Therefore, by Ito’s formula, under $\mathds{P}$, we have
(33) $dw(Y_{t})=Lw(Y_{t})\,\mathrm{d}t+2\epsilon\nabla\log\tilde{q}(Y_{t})\cdot\nabla w(Y_{t})\,\mathrm{d}t+\sqrt{2\epsilon}\nabla w(Y_{t})\cdot\mathrm{d}W_{t},\\$
where $W_{t}$ is the $\mathds{P}$-Brownian motion defined in Theorem 3.11 Now let
$\tilde{v}:=-2\epsilon\log(\tilde{q})\text{ and }v:=-2\epsilon T.$
By a calculation similar to the proof of Lemma 3.6, we have
$\frac{L\tilde{q}}{\tilde{q}}=-\frac{1}{2\epsilon}\left(L\tilde{v}-\frac{1}{2}|\nabla\tilde{v}|^{2}\right)\text{ and }\frac{LT}{T}=-\frac{1}{2\epsilon}\left(Lv-\frac{1}{2}|\nabla v|^{2}\right).$
Also, $w=-\frac{1}{2\epsilon}(\tilde{v}-v)$. Plugging these expressions into  (33), we get
$\mathrm{d}w(Y_{t})$ $=\left(-\frac{1}{2\epsilon}L\tilde{v}(Y_{t})+\frac{1}{2\epsilon}Lv(Y_{t})+\frac{1}{2\epsilon}\nabla\tilde{v}(Y_{t})\cdot\nabla(\tilde{v}-v)(Y_{t})\right)\,\mathrm{d}t$
$\qquad+\sqrt{2\epsilon}\nabla w(Y_{t})\cdot\mathrm{d}B_{t}$
$=-\frac{1}{2\epsilon}\left(L\tilde{v}(Y_{t})-\frac{1}{2}|\nabla\tilde{v}(Y_{t})|^{2}\right)\mathrm{d}t$
$\qquad+\frac{1}{2\epsilon}\left(Lv(Y_{t})-\frac{1}{2}|\nabla v(Y_{t})|^{2}\right)\mathrm{d}t$
$\qquad+\frac{1}{2\epsilon}\left(\frac{1}{2}|\nabla\tilde{v}(X_{2})|^{2}+\frac{1}{2}|\nabla v(X_{2})|^{2}-\nabla\tilde{v}(Y_{t})\cdot\nabla v(Y_{t})\right)\mathrm{d}t$
$\qquad+\sqrt{2\epsilon}\nabla w(Y_{t})\cdot\mathrm{d}B_{t}$
$=\left(\frac{L\tilde{q}}{\tilde{q}}(Y_{t})-\frac{LT}{T}(Y_{t})\right)\mathrm{d}t$
$\qquad+\epsilon|\nabla w(Y_{t})|^{2}\mathrm{d}t+\sqrt{2\epsilon}\nabla w(Y_{t})\cdot\mathrm{d}W_{t},$
and the result follows.
We use Lemma 5.1 to show that our estimator of relative entropy differences has finite variance, at least under certain conditions on $\tilde{q}$.
Lemma B.2.
Assume that $\mathds{E}_{\mathds{P}}[\tau^{2}]<\infty$, that for some $C>0$,
[…]
for all $x\in\bar{\mathscr{T}}$, and that for some $M>0$,
$\frac{1}{M}\leq\lvert m(x)\rvert\leq M\text{ and }\frac{1}{M}\leq\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)\leq M$
for all $x\in\partial A$. If so,
$\operatorname{var}_{\mathds{P}}\left(\log\left(\frac{m(\tilde{Y}^{1}_{0})}{\lvert\nabla\tilde{q}(\tilde{Y}^{1}_{0})\rvert\exp(-U(\tilde{Y}^{1}_{0})/\varepsilon)}\right)+\log\tilde{q}(\tilde{Y}^{1}_{\tau})-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(\tilde{Y}^{1}_{s})\,\mathrm{d}s\right)<\infty,$
so $\lim_{\tilde{N}\rightarrow\infty}\tilde{I}_{\tilde{N}}=\tilde{I}$ in $L^{2}(\mathds{P})$.
Proof B.3.
The crux of the proof is to show that
$\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)^{2}\right]<\infty$
even though $\frac{L\tilde{q}}{\tilde{q}}$ may be singular on $\partial A$. Taking $T=q$ in Lemma 5.1 yields
$\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s$ $=\log\frac{r(Y_{\tau})}{r(Y_{0})}-\varepsilon\int_{0}^{\tau}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s-\sqrt{2\varepsilon}\int_{0}^{\tau}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}$
where $r=\tilde{q}/q$ and $W_{s}$ is the $\mathds{P}$-Brownian motion in Theorem 3.11. Therefore,
$\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)^{2}\right]^{\frac{1}{2}}$ $\leq\mathds{E}_{\mathds{P}}\left[\left(\log\frac{r(Y_{\tau})}{r(Y_{0})}\right)^{2}\right]^{\frac{1}{2}}+\varepsilon\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right)^{2}\right]^{\frac{1}{2}}$
$\qquad\qquad\qquad+\sqrt{2\varepsilon}\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]^{\frac{1}{2}}$
(34) $\leq 2C+\varepsilon C\mathds{E}_{\mathds{P}}[\tau^{2}]+\sqrt{2\varepsilon}\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{\tau}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]^{\frac{1}{2}}.$
We bound the third term in (34) using the Ito isometry, which entails some minor but tiresome technical difficulties, since standard forms of the isometry only apply to integrals up to finite, deterministic times. Define
$S_{t}=\int_{0}^{t}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}.$
We will approximate $S_{\tau}$ by $S_{\tau\wedge n}$ for $n\in\mathbb{N}$ to apply the Ito isometry; we claim that $S_{\tau\wedge n}$ is a Cauchy sequence in $L^{2}(\mathds{P})$. To see this, let $n<n^{\prime}\in\mathbb{N}$, and observe that
$\mathds{E}_{\mathds{P}}[(S_{\tau\wedge n^{\prime}}-S_{\tau\wedge n})^{2}]$ $=\mathds{E}_{\mathds{P}}\left[\left(\int_{\tau\wedge n}^{\tau\wedge n^{\prime}}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]$
$=\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{n^{\prime}}\mathds{1}_{\tau\wedge n\leq s<\tau\wedge n^{\prime}}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]$
$=\mathds{E}_{\mathds{P}}\left[\int_{0}^{n^{\prime}}\mathds{1}_{\tau\wedge n\leq s<\tau\wedge n^{\prime}}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right]$
$\leq\mathds{E}_{\mathds{P}}\left[\int_{0}^{n^{\prime}}\mathds{1}_{\tau\geq n}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right]$
$\leq C^{2}\mathds{P}[\tau\geq n].$
The third equality above follows from the Ito isometry, since $\tau$ is a stopping time, hence $\mathds{1}_{\tau\wedge n\leq s<\tau\wedge n^{\prime}}$ is adapted. The first inequality holds, since $\tau<n$ implies $\mathds{1}_{\tau\wedge n\leq s<\tau\wedge n^{\prime}}=0$. We assume that $\mathds{P}[\tau<\infty]=1$, so $\lim_{n\rightarrow\infty}\mathds{P}[\tau\geq n]=0$, hence $S_{\tau\wedge n}$ is Cauchy. In fact, we have $\lim_{n\rightarrow\infty}S_{\tau\wedge n}=S_{\tau}$ in $L^{2}(\mathds{P})$, since $\mathds{P}[\tau<\infty]=1$ implies that $S_{\tau\wedge n}$ converges to $S_{\tau}$ almost surely. We conclude that
$\mathds{E}_{\mathds{P}}[S_{\tau}^{2}]$ $=\lim_{n\rightarrow\infty}\mathds{E}_{\mathds{P}}[S_{\tau\wedge n}^{2}]$
$=\lim_{n\rightarrow\infty}\mathds{E}_{\mathds{P}}\left[\left(\int_{0}^{n}\mathds{1}_{s\leq\tau\wedge n}\nabla\log r(Y_{s})\cdot\mathrm{d}W_{s}\right)^{2}\right]$
$=\lim_{n\rightarrow\infty}\mathds{E}_{\mathds{P}}\left[\int_{0}^{n}\mathds{1}_{s\leq\tau\wedge n}\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right]$
$\leq\lim_{n\rightarrow\infty}C^{2}\mathds{E}_{\mathds{P}}[\tau\wedge n]$
$=C^{2}\mathds{E}_{\mathds{P}}[\tau].$
It follows that
$\operatorname{var}_{\mathds{P}}\left(\log\left(\frac{m(\tilde{Y}^{1}_{0})}{\lvert\nabla\tilde{q}(\tilde{Y}^{1}_{0})\rvert\exp(-U(\tilde{Y}^{1}_{0})/\varepsilon)}\right)+\log\tilde{q}(\tilde{Y}^{1}_{\tau})-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(\tilde{Y}^{1}_{s})\,\mathrm{d}s\right)<\infty.$
We leave the remaining details to the reader.
We now prove Theorem 4.5 on the differentiability of $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ in $\theta$.
Proof B.4 (Proof of Theorem 4.5).
Since we assume that $\frac{Lq_{\theta}}{q_{\theta}}$ and $\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}$ are bounded, for each $\omega\in\Omega$,
$\nabla_{\theta}\int_{0}^{\tau(\omega)}\frac{Lq_{\theta}}{q_{\theta}}(X_{s}(\omega))\,\mathrm{d}s=\int_{0}^{\tau(\omega)}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(X_{s}(\omega))\,\mathrm{d}s.$
Moreover, since we assume $\eta_{\theta}\geq\frac{1}{M}>0$, and since $q_{\theta}$ is differentiable, $\log\eta_{\theta}$ is differentiable. Therefore,
(35) $\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}$ $=\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\left(\frac{\nabla_{\theta}q_{\theta}(X_{\tau})}{q_{\theta}(X_{\tau})}-\frac{\nabla_{\theta}\eta_{\theta}}{\eta_{\theta}}-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s\right).$
One now wants to interchange differentiation with respect to $\theta$ and expectation over $\mathds{Q}$ to derive a formula for $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$. Write
$g(x)=x\log x$
so that
$D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})=\mathds{E}_{\mathds{Q}}\left[g\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right].$
We claim that
(36) $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})=\nabla_{\theta}\mathds{E}_{\mathds{Q}}\left[g\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right]=\mathds{E}_{\mathds{Q}}\left[g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right].$
Assumption 4.3 does not imply that $\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}$ is bounded uniformly in $\theta$ by a function in $L^{1}(\mathds{Q})$, so textbook results related to differentiating under the integral do not apply. To verify (36), we return to the definition of the derivative. Given any $h,\theta\in\mathbb{R}^{k}$, by the mean value theorem, for some $s:\Omega\rightarrow[0,1]$, we have
$\frac{1}{\lVert h\rVert}\mathds{E}_{\mathds{Q}}$ $\left[g\left(\frac{\mathrm{d}\mathds{P}_{\theta+h}}{\mathrm{d}\mathds{Q}}\right)-g\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)-g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot h\right]$
$=\mathds{E}_{\mathds{Q}}\left[g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}-g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right]$
$=\mathds{E}_{\mathds{Q}}\left[\left\{g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)-g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+h}}{\mathrm{d}\mathds{Q}}\right)\right\}\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right$.
$\qquad\qquad+\left.g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)\left\{\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}-\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right\}\cdot\frac{h}{\lVert h\rVert}\right]$
$=\mathds{E}_{\mathds{P}_{\theta}}\left[\left\{\log\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)-\log\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\right\}\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right$.
$\qquad\qquad+\left\{\log\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)+1\right\}\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\frac{\mathrm{d}\mathds{Q}}{\mathrm{d}\mathds{P}_{\theta}}$
$\qquad\qquad-\left.\left\{\log\left(\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\right)+1\right\}\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\cdot\frac{h}{\lVert h\rVert}\right]$
$=:\mathds{E}_{\mathds{P}_{\theta}}[R(\theta;h)].$
As $h\rightarrow 0$, the integrand $R(\theta;h)$ in the expectation with respect to $\mathds{P}_{\theta}$ above converges pointwise to zero. We now show that $R(\theta;h)$ is dominated in $L^{1}(\mathds{P}_{\theta})$, uniformly in $h$ for sufficiently small $h$. If so, $\lim_{h\rightarrow 0}\mathds{E}_{\mathds{P}_{\theta}}[R(\theta;h)]=0$, which verifies differentiability of $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$ and (36). By Assumption 4.3, we have
$\left\lvert\frac{\mathrm{d}\mathds{P}_{\theta+sh}}{\mathrm{d}\mathds{Q}}\frac{\mathrm{d}\mathds{Q}}{\mathrm{d}\mathds{P}_{\theta}}\right\rvert$ $\leq\frac{\eta_{\theta}}{\eta_{\theta+sh}}\exp\left(\int_{0}^{\tau}\left\lvert\frac{Lq_{\theta}}{q_{\theta}}(Y_{t})-\frac{Lq_{\theta+sh}}{q_{\theta+sh}}(Y_{t})\right\rvert\,\mathrm{d}t\right)$
$\leq M^{2}\exp(C\lvert h\rvert\tau),$
[…]
$\leq\lvert\log\zeta\rvert+\lvert\log M\rvert+C\tau,$
and
[…]
$\leq M^{2}+C\tau.$
Therefore, when $\lvert h\rvert\leq\frac{\gamma(\theta)}{2C}$,
$\lvert R(\theta;h)\rvert$ $\leq 2(\lvert\log\zeta\rvert+\lvert\log M\rvert+C\tau)(M^{2}+C\tau)$
$\qquad+(1+\lvert\log\zeta\rvert+\lvert\log M\rvert+C\tau)(M^{2}+C\tau)$
$\qquad+(1+\lvert\log\zeta\rvert+\lvert\log M\rvert+C\tau)(M^{2}+C\tau)M^{2}\exp\left(\frac{\gamma(\theta)}{2}\tau\right).$
Under Assumption 4.3, the random variable on the right-hand-side above is in $L^{1}(\mathds{P}_{\theta})$, which concludes the proof of differentiability.
We now derive a more convenient formula for the derivative. As a first step, we observe that one can again exchange expectation and differentiation to obtain
$0$ $=\nabla_{\theta}\mathds{E}_{\mathds{P}_{\theta}}[\Omega]$
$=\nabla_{\theta}\mathds{E}_{\mathds{Q}}\left[\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{E}_{\mathds{Q}}\left[\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
(37) $=\mathds{E}_{\mathds{P}_{\theta}}\left[\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right].$
Now we have
$\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}|\mathds{Q})$ $=\mathds{E}_{\mathds{Q}}\left[g^{\prime}\left(\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{E}_{\mathds{Q}}\left[\left(\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}+1\right)\nabla_{\theta}\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\mathds{E}_{\mathds{P}_{\theta}}\left[\left(\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}+1\right)\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right]$
$=\operatorname{cov}_{\mathds{P}_{\theta}}\left(\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}},\nabla_{\theta}\log\frac{\mathrm{d}\mathds{P}_{\theta}}{\mathrm{d}\mathds{Q}}\right)$
$=\operatorname{cov}_{\mathds{P}_{\theta}}\left(\log\frac{\zeta}{\eta_{\theta}}+\log q_{\theta}(X_{\tau})-\int_{0}^{\tau}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s,\right$.
$\qquad\qquad\quad\left.\frac{\nabla_{\theta}q_{\theta}(X_{\tau})}{q_{\theta}(X_{\tau})}-\frac{\nabla_{\theta}\eta_{\theta}}{\eta_{\theta}}-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s\right)$
$=\operatorname{cov}_{\mathds{P}_{\theta}}\left(\log q_{\theta}(X_{\tau})-\int_{0}^{\tau}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s,\right$.
$\qquad\qquad\quad\left.\frac{\nabla_{\theta}q_{\theta}(X_{\tau})}{q_{\theta}(X_{\tau})}-\int_{0}^{\tau}\nabla_{\theta}\frac{Lq_{\theta}}{q_{\theta}}(X_{s})\,\mathrm{d}s\right)$
The fourth equality above follows from (37), and the last follows since $\log\frac{\zeta}{\eta_{\theta}}$ and $\frac{\nabla_{\theta}\eta_{\theta}}{\eta_{\theta}}$ are constants that do not depend on $\omega$.
Appendix C Detailed Description of Numerical Experiments
Here, we explain the numerical experiments outlined in Section 6 in more detail.
C.1 Details related to Section 6.2
We used FENICS to compute the committor function by a finite element method [3]. We minimized the Ritz form (25) with the prescribed boundary conditions over a P1 Lagrange finite element space based on a Delaunay triangulation of a regular rectangular grid of size $256\times 1024$ covering the domain $a,b]\times[-4,4]$.
C.2 Details related to Section 6.4
When calculating an approximate committor by gradient descent, we took as the function $w_{\theta}(x)$ a $4\times 16$ tensor product of cardinal cubic B-splines covering $a,b]\times[-3,3]$. That is, for $\theta\in\mathbb{R}^{4\times 16}$ and $x\in\mathbb{R}^{2}$, we defined
$w_{\theta}(x)=\sum_{i=0}^{3}\sum_{j=0}^{15}\theta_{ij}B_{3}\left(\frac{x_{1}-a}{h_{1}}-i\right)B_{3}\left(\frac{x_{2}-(-3)}{h_{2}}-j\right),$
where
$h_{1}=\frac{b-a}{4-1},\text{ }h_{2}=\frac{3-(-3)}{16-1}$
and $B_{3}$ is the cardinal cubic B-spline
$B_{3}(x):=\frac{1}{6}\begin{cases}(2-\lvert x\rvert)^{3}&\text{ if }1\leq\lvert x\rvert\leq 2\\ 4-6x^{2}-3\lvert x\rvert^{3}&\text{if }\lvert x\rvert\leq 1\\ 0&\text{otherwise}\end{cases}$
For the Adam optimizer, we used the standard parameters listed in [19], except that we took $\epsilon=10^{-4}$ instead of $10^{-8}$ to mitigate Adam’s tendency to make wild changes to some entries of $\theta$ when trajectories visit rare regions far from the transition tube. We took an exponentially decreasing step size schedule
$\eta_{n}=0.1^{1+n/512}.$
Although it is well known that stochastic optimization methods may not converge to a minimizer for such a rapidly decreasing schedule, we obtained better results with an exponentially decreasing schedule than with polynomially decreasing schedules for which convergence is guaranteed. Our batch size was $64$. That is, we calculated $64$ trajectories to estimate $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{P}_{\theta}|\mathds{Q})$ for each Adam step.
To simplify the sampling of initial points of trajectories, we approximated the reactive flux distribution $\tilde{m}$ corresponding to $\tilde{q}$ by a discrete distribution
$m_{d}:=\frac{1}{\eta_{d}}\sum_{i=1}^{N_{d}}\lvert\nabla\tilde{q}(x_{i})\rvert\exp(-U(x_{i})/\varepsilon)\delta_{x_{i}}$
supported on the grid
$x_{i}=\left(a,-3+\frac{6}{N_{d}-1}\right)\text{ for }0\leq i\leq N_{d}-1$
covering the interval $\{a\}\times[-3,3]\subset\partial A$ with $N_{d}=1024$. Here,
$\eta_{d}=\sum_{i=1}^{N_{d}}\lvert\nabla\tilde{q}(x_{i})\rvert\exp(-U(x_{i})/\varepsilon)$
is a normalizing constant.
Finally, for efficiency, instead of computing $q_{1}$ defined in (24) by quadrature at each point of every trajectory biased by an approximate comittor of the form $q_{1}(x_{1})\exp(w_{\theta}(x)(b-x_{1}))$ as in (31), we interpolated a quadrature estimate of $q_{1}$ by a B-spline plus a linear function on a uniform grid of $16$ points covering $[a,b]$. We verified the interpolating spline to be correct to approximately five significant figures.
C.3 Details related to Section 6.5
We used the delta method to produce the standard errors in Tables 1 and 2. That is, to estimate the asymptotic variance of the self-normalized importance sampling estimator
$\bar{\tau}:=\frac{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)\tau_{k}}{\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)},$
of the expected crossover time $\mathds{E}_{\mathds{Q}}[\tau]$, we first calculate the sample averages
$\tilde{\tau}$ $:=\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)\tau_{k},$
$\bar{Z}$ $:=\frac{1}{N}\sum_{k=1}^{N}\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right),$
and the covariance matrix
$\bar{\Sigma}:=\begin{pmatrix}\Sigma_{\tilde{\tau},\tilde{\tau}}&\Sigma_{\tilde{\tau},\bar{Z}}\\ \Sigma_{\bar{Z},\tilde{\tau}}&\Sigma_{\bar{Z},\bar{Z}}\\ \end{pmatrix}$,
where $\Sigma_{\tilde{\tau},\bar{Z}}$ is the empirical covariance
$\Sigma_{\tilde{\tau},\bar{Z}}:=\frac{1}{N-1}\sum_{k=1}^{N}\left\{\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)\tau_{k}-\tilde{\tau}\right\}\left\{\exp\left(\int_{0}^{\tau_{k}}\frac{L\tilde{q}}{\tilde{q}}(Y^{k}_{s})\,\mathrm{d}s\right)-\bar{Z}\right\},$
of $\tilde{\tau}$ with $\bar{Z}$ and the other entries of $\bar{\Sigma}$ are defined similarly. We then define $g(x,y)=x/y$ so that $\bar{\tau}=g(\tilde{\tau},\bar{Z})$, and the delta method [5] yields
$s^{2}:=\frac{1}{N}\nabla g(\tilde{\tau},\bar{Z})^{\text{t}}\bar{\Sigma}\nabla g(\tilde{\tau},\bar{Z})=\frac{1}{N}\begin{pmatrix}1/\bar{Z},&-\tilde{\tau}/\bar{Z}^{2}\end{pmatrix}\bar{\Sigma}\begin{pmatrix}1/\bar{Z}\\ -\tilde{\tau}/\bar{Z}^{2}\end{pmatrix}$.
as an estimate of the asymptotic variance of $\bar{\tau}$.